### v1

# Can Large Language Models Act as Ensembler for Multi-GNNs?

Abstract.
Graph Neural Networks (GNNs) have emerged as powerful models for learning from graph-structured data. However, GNNs lack the inherent semantic understanding capability of rich textual nodes attributes, limiting their effectiveness in applications. On the other hand, we empirically observe that for existing GNN models, no one can consistently outperforms others across diverse datasets. In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. The model first aligns multiple GNNs, mapping the representations of different GNNs into the same space. Then, through LoRA fine-tuning, it aligns the space between the GNN and the LLM, injecting graph tokens and textual information into LLMs. This allows LensGNN to ensemble multiple GNNs and leverage LLM’s strengths, leading to a deeper understanding of both textual semantic information and graph structural information. Experimental results show that LensGNN  outperforms existing models. This research advances text-attributed graph ensemble learning by providing a robust, superior solution for integrating semantic and structural information. We provide our code and data here: https://anonymous.4open.science/r/EnsemGNN-E267/.
Large Language Models, Text-attributed Graph, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. INTRODUCTION
Graphs are structured data that captures the inter-relations between entities in the real world. To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced (Kipf and Welling, 2016) and message passing is the main mechanism for neighborhood aggregation (Hamilton et al., 2017a; Veličković et al., 2017). GNNs have demonstrated significant success across various applications such as social network analysis (Hamilton et al., 2017a), recommendation systems (Wang et al., 2019), and molecular property prediction (Gilmer et al., 2017).
Despite the success, two major challenges remain unresolved. First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes (Zhao et al., 2019). This can lead to the loss of valuable semantic information during the learning process, limiting the effectiveness of GNNs in applications where node features contain meaningful texts, such as citation networks and social media platforms. Second, while there have been a series of GNNs proposed, no one has been shown to consistently lead others across datasets of various domains. For example, we compare the node classification performance of four representative GNNs, namely, APPNP (Gasteiger et al., 2018), GAT (Veličković et al., 2017), GCN (Kipf and Welling, 2016) and GIN (Xu et al., 2019) on three benchmark datasets (Yan et al., 2023): Cora, Citeseer and PubMed. For fairness, we set the equal number of hidden representation layers and also the same dimensionality size. The results given in Figure 1 demonstrate that the winner varies across the datasets. The challenge of selecting the optimal GNN for a given dataset remains unresolved, as different GNN architectures exhibit varying strengths. This also restricts the wide applicability of GNNs.
[…]
Comparison of node classification accuracy for various GNN models across different datasets. To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
Figure 1. Comparison of node classification accuracy for various GNN models across different datasets. To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
To remedy the incapability of GNNs in semantic understanding, existing works (Qin et al., 2024; Li et al., 2024; Wang et al., 2024; Ye et al., 2024; Fang et al., 2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation. The inherent advantage of GNNs in utilizing graph structure, fused with the strength of LLMs in semantic understanding, mutually enhances both models and leads to superior performance in downstream tasks. However, how to select the optimal GNN in different scenarios remains a gap. Since the effectiveness of GNNs varies across datasets, there naturally arises a question: Can we develop an ensembled model for multi-GNNs? The ensembled model is expected to integrate the strengths of multiple GNNs and can consistently perform well across datasets. To further leverage the power of LLMs in text understanding, we thus upgrade the question: Can LLMs act as ensembler for multi-GNNs?
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. The model captures semantic information from node textual attributes using LLMs, while simultaneously enabling LLMs to implicitly evaluate and combine the strengths of multiple GNNs. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
  • We propose a novel method LensGNN to ensemble multi-GNNs with LLMs. To our best knowledge, we are the first work to attempt the trial. LensGNN combines the strengths of both GNNs and LLMs, which can be used to avoid the intricate process of manual GNN selection on various datasets.
  • We introduce an effective graph prompt tuning paradigm, which enhances LLMs’ capability in comprehending graph topology with two phases of alignment. The multi-GNN alignment maps multi-GNNs into the same low-dimensional space while the GNN-LLM alignment injects graph token representations into LLMs.
  • We conduct extensive experiments to show the effectiveness of LensGNN. In particular, we compare LensGNN with 17 state-of-the-art competitors. Our experimental results show that LensGNN outperforms baselines in a majority of cases, demonstrating its superior performance.
2. RELATED WORK
GNNs (Kipf and Welling, 2016; Xu et al., 2019; Veličković et al., 2017; Hamilton et al., 2017b) have been widely used in learning from graph-structured data. Existing GNN studies have been conducted on various types of graphs such as heterophilic graphs (Zhu et al., 2020; Pei et al., 2020; Bi et al., 2024), signed graphs (Derr et al., 2018; Huang et al., 2019) and temporal graphs (Rossi et al., 2020; Longa et al., 2023; Xu et al., 2020). Further, to improve the effectiveness of GNNs, there are also works (Nguyen et al., 2023; Rusch et al., 2023; Black et al., 2023) devoted to solving the notorious over-smoothing (Chen et al., 2020) and over-quashing problems (Akansha, 2023). In parallel, to scale GNNs to large-scale graphs, some works (Liao et al., 2022) focus on enhancing the model efficiency. A comprehensive GNN-related survey can be found at (Wu et al., 2020). Despite their advantages, GNNs struggle to process textual attributes effectively, as traditional GNNs lack the semantic understanding capability that is necessary to handle text-attributed graphs (TAGs). In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) (Zoph and Le, 2017). However, it is very computationally expensive, which necessitates new exploration.
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies (Jin et al., 2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, where models like GraphLLM (Chai et al., 2023) utilizes few-shot learning to address graph reasoning challenges. ENGINE (Zhu et al., 2024) is designed to efficiently integrate textual and topological information by adding a tunable lightweight GNN side structure alongside each layer of LLMs.Similarly, heuristic reasoning approaches, including Chain-of-Thought (CoT) (Wei et al., 2022) and models like StructGPT (Jiang et al., 2023), have demonstrated improved graph-based inference capabilities. Another notable category is to employ LLMs as encoders, as seen in the works of TextGNN (Zhu et al., 2021) and AdsGNN (Li et al., 2021), which focus on optimization strategies for encoding graph data. Employing LLMs as encoders involves using large language models to extract textual features, which are then used as initial feature vectors for nodes or edges in a graph. These feature vectors are passed into graph neural networks (GNNs) for further processing, allowing the model to incorporate both the textual information from the LLM and the structural information from the graph. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods such as LTRN (Zhang et al., 2021) and GLEM (Zhao et al., 2022), where iterative training helps improve cross-modality learning. LLMs and GNNs are aligned through iterative pseudo-label training (prediction alignment) or by using contrastive learning to align their latent representations (latent space alignment). Finally, graph-empowered LLMs like GreaseLM (Zhang et al., 2022) and DRAGON (Yasunaga et al., 2022) illustrate the potential of enhancing language models with graph-based knowledge to enrich semantic understanding. They integrate the strengths of large language models and graph structures by modifying Transformer architectures to encode text and graph data simultaneously, allowing for a more comprehensive understanding of the interconnected information within graphs. These diverse approaches highlight the evolving landscape of LLM and GNN integration. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
3. PRELIMINARIES
Text-attributed graph. A text-attributed graph is a graph where nodes and edges have textual attributes. In this work, we only consider the case where only nodes have textual attributes. Formally, a text-attributed graph can be defined as $G=(V,E,X^{(v)})$, where $V=\{v_{1},v_{2},\ldots,v_{|V|}\}$ represents the set of nodes, and $E=\{e_{1},e_{2},\ldots,e_{|E|}\}$ represents the set of edges, with $|V|=N$ indicating the total number of nodes in the graph. $X^{(v)}=\{x^{(v)}_{1},x^{(v)}_{2},\ldots,x^{(v)}_{|V|}\}$ denote the attributes on nodes, respectively, which are strings. It can be represented as $G=(V,E,\{x_{n}\}_{n\in V})$. Additionally, we define $A$ as the adjacency matrix of graph $G$, where the matrix size is $N\times N$. In an unweighted graph, $A(i,j)=1$ indicates that there is an edge between nodes $v_{i}$ and $v_{j}$, while $A(i,j)=0$ indicates that there is no edge between them.
Graph Neural Networks. Traditional GNNs are a class of deep learning models designed for handling graph-structured data. The basic architecture of a GNN includes a node representation layer and a message-passing layer. In the node representation layer, each node $v$ is assigned a feature vector $x_{v}$. In the message-passing layer, the representation vector $h_{v}^{(t)}$ of node $v$ is updated after the $t$-th iteration using the following formula:
(1) $h_{v}^{(t)}=\text{UPDATE}\left(\text{AGGREGATE}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$
where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$. The AGGREGATE function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state. Through this iterative process, GNNs are able to learn increasingly rich node representations, capturing both local and global structural information in the graph. The specific implementations of the AGGREGATE and UPDATE functions can vary depending on the particular GNN model.
4. METHODOLOGY
To enable the LLM to integrate multiple GNNs, alignment is primarily required in two aspects. First, the output representations of all the GNNs need to be aligned into a common vector space. Second, there must be alignment between the output representations of the GNNs and the semantic space required by the LLM. Based on this idea, we have designed the model in the following two steps. The structure of the whole model is shown in Fig. 2.
4.1. Aligning multi-GNNs
Given multiple GNNs, they could generate node representations in different low-dimensional spaces. Therefore, before ensembling GNNs, we need to first align them.
Figure 2. The overall process of LensGNN. Step 1 is the workflow of the pre-training phase of LensGNN’s GNNs. We employ a language pre-training encoder to extract semantic information from the textual attributes of nodes. Concurrently, we train the GNN encoder to capture structural information. During this phase, a MLP is utilized as a classifier, and the model is trained through gradient feedback. Step 2 is the workflow of the LoRA (Low-Rank Adaptation) supervised fine-tuning stage of LensGNN’s LLM. At this stage, we freeze the parameters of the GNNs and infer the hidden representations of the graph nodes. We segment GNN representations and convert them into graph tokens that LLM can read. The graph tokens, node text tokens, and other prompts are concatenated to form an instruction, which is then input into the LLMs for LoRA supervised fine-tuning.
[…]
In the pre-training phase, LensGNN consists of the following architecture: a pre-trained language encoder with frozen parameters, a GNN layer, a linear mapping layer, and an MLP classifier, all of which have parameters that are updated during training.In the fine-tuning phase of LensGNN, the parameters of the pre-trained language encoder and the GNN layers are frozen, and only the Lora parameters of the LLM are updated. In addition, the size of Hidden Representation is (1, t * e), where e is the length of the enbedding of each token in LLM, and t is the number of tokens each node representation is converted into.
Node feature initialization. To utilize the rich semantic information of textual features, we use pre-trained language models to initialize them. Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT (Reimers, 2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
(2) $\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$
This process allows texts of any dataset and length to be converted into the same feature space.
GNN representations. Next, the model takes the text representations obtained from Eq. 2 for each node and feeds them into multiple GNNs. For the $k$-th GNN, we denote the node embedding matrix $H^{[k]}$ generated by the final layer as:
(3) $H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$
where $\tilde{X}$ is the initial node embedding matrix from Eq. 2 and $\Theta_{G_{k}}$ is the learnable parameters of the $k$-th GNN. For each node $v_{i}$, let $h_{i}^{[k]}$ be its embedding vector, which is the $i$-th row in $H^{[k]}$.
Reshape node representations. The output of each GNN is then connected to a linear layer, which transforms node representations to the dimensionality of hidden embeddings in LLMs. This step paves the way for the subsequent GNN and LLM alignment. Details will be given in the next section. For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
(4) $\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$
Note that the shape of $\tilde{h}^{[k]}_{i}$ is a one-dimensional vector of length $t\times e$, where $t$ is a hyperparameter indicating the number of graph tokens each node representation is mapped to. The dimensionality $e$ comes from the hidden embedding layer of the LLM used, where each token is mapped to an embedding of shape $(1,e)$.
Multi-GNN alignment. To align node representations from multiple GNNs, we next feed them into a shared linear layer, which serves as the classifier. During the training time, we alternatively input representations from different GNNs into the classifier and train them sequentially. This training approach allows node representations from multi-GNNs to be aligned, which facilitates the integration of multi-GNNs. After training, the parameters of GNNs and the linear layer in Eq. 4 will be frozen, which are then used in aligning GNNs and LLMs. The overall procedure for GNN alignment is summarized in the top of Fig. 2.
In summary, during the multi-GNN alignment step, we train each GNN to ensure that their outputs reside in the same vector space. By incorporating the node feature initialization step that utilizes language pre-trained models, our model gains the ability to extract semantic and structural information from the nodes in the text graph. Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the integration of extensive graph structural information into the LLM in the subsequent steps.
4.2. Ensembling multi-GNNs with LLM
After GNNs are trained, we can get node representations capturing rich graph structural information. Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure (Guo et al., 2023). Therefore, in the second stage, we empower LLMs to comprehend graph representations. Further, with specially designed prompts, we enable these models to implicitly perform ensembling.
Align GNNs and LLMs. The key step in enabling LLMs to understand GNN tokens lies in aligning GNN representations, which encapsulate rich structural information, with the semantic space required by LLMs. This process necessitates fine-tuning LLMs. However, due to the extensive number of parameters in these models, fine-tuning requires significant computational resources, which is challenging for practical applications. As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector (Tang et al., 2024), mapping GNN representations into the semantic space. While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
[…]
Figure 3. A prompt template. The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific node text inputs and their multiple GNN representations.
Prompt design and Multi-GNN ensembling. After aligning graph structure and semantics, LLM can understand graph representations. We next design prompts and take LLM as the ensembler to integrate multi-GNNs. Our goal is to leverage the strengths of both multi-GNNs and LLM to output more accurate predictions. The prompt design follows three key principles:
  1. (1)
    Simultaneous input of textual strings from a node and its adjacent neighbors, graph tokens of the node, and task specification. Existing studies (Kipf and Welling, 2016; Veličković et al., 2017; Verma et al., 2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction. Further, clear task instructions are necessary for accurate prediction.
  2. (2)
    Differentiation between text tokens and graph tokens. It directs LLM to correctly distinguish between textual tokens and graph tokens within different segments of the prompt, thus preventing confusion.
  3. (3)
    Guidance for learning from multi-GNNs: It leads LLM to implicitly learn how to effectively combine strengths of different GNNs.
Based on the above principles, we provide a prompt template used in our experiments, as shown in Figure 3. Details on the prompt are given in the caption of the figure. In practice, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
5. EXPERIMENT
In this section, we conduct extensive experiments to evaluate the performance of our model across various datasets, aiming to answer the following questions:
  • RQ1: How does LensGNN perform compared to SOTA baseline models?
  • RQ2: What is the importance of different components in LensGNN?
  • RQ3: Can small language models serve as multi-GNN ensembler?
  • RQ4: How efficient is our method LensGNN ?
  • RQ5: Is our method LensGNN insensitive to model hyper-parameters?
Table 1. Datasets statistics.
Dataset # Nodes # Edges # Classes
Cora 2,708 5,429 7
PubMed 19,717 44,338 3
ogbn-arXiv 169,343 1,166,243 40
Table 2. Hyperparameter Search Spaces.
Hyperparameter Search space
learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
dropout $[0,0.25]$
loraplus lr ratio $\{16.0,24.0,32.0\}$
Graph token $t$$\{1,2,4,8,16\}$
Table 3. The classification performance over the methods on 3 datasets. We highlight the best score on each dataset in bold and underline the runner-up’s. The - means that we did not obtain results from their original paper.
Model type Model Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
MLP MLP 66.42 82.40 61.47
GNN GCN 85.97 83.78 70.81
GAT 86.71 83.59 70.05
GIN 85.60 82.03 67.03
GraphSAGE 84.87 87.01 70.35
Graphormer 80.41 88.75 72.81
GT 86.42 88.24 -
CoarFormer 88.69 89.75 71.66
Language Model BERT 80.15 93.91 72.78
SentenceBERT 78.82 92.49 71.42
DeBERTa 77.79 93.45 72.90
LLM Based DGTL 81.10 87.10 -
SNS-GPT4 82.50 93.80 74.40
InstructGraph-INS 89.33 81.09 -
InstructGLM 90.77 94.62 75.70
GAugLLM - 83.68 74.15
Baichuan2-13B 13.65 36.04 4.79
LensGNN-[GCN+GAT] 88.56 94.11 75.78
LensGNN-[GCN+GIN] 91.88 94.47 75.89
LensGNN-[GAT+GIN] 88.19 95.43 74.31
LensGNN-ALL 90.40 95.68 75.91
5.1. Experimental Settings
Datasets. We use three benchmark datasets whose details are given as follows. Statistics of these datasets are summarized in Table 1.
Cora is a citation network containing 2,708 research papers. These papers are categorized into seven classes. In this graph, each node represents a scientific paper, and each edge represents a citation. Bag-of-words representation is used as the feature vector for each node, which is also assigned a label indicating its research field. The task for this dataset is node classification.
PubMed is a citation network dataset containing 19,717 research papers from the biomedical field. These papers are categorized into three classes, and each paper is represented by a bag-of-words feature vector. In this graph, each node represents a scientific paper, and each edge represents a citation. The task associated with this dataset is node classification.
ogbn-arXiv is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv. In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships. These papers are categorized into 40 subject areas. We construct node attributes using the title and abstract of the papers. The task associated with this dataset is node classification.
Baselines. We compare LensGNN with 17 baselines, which can be categorized into four types. The first type involves (1) MLP. The second type involves (2) GCN (Kipf and Welling, 2016), (3) GAT (Veličković et al., 2017), (4) GIN (Xu et al., 2019), (5) GraphSAGE (Hamilton et al., 2017a), (6) Graphormer (Ying et al., 2021), (7) GT (Dwivedi and Bresson, 2012) and (8) CoarFormer (Kuang et al., 2024) seven popular GNNs for graph tasks. For fairness, we use the node representations obtained from a pre-trained LM as input to the GNN models for learning node classification. The third type involves (9) BERT (Devlin, 2018), (10) SentenceBERT (Reimers, 2019) and (11) DeBERTa (He et al., 2020) three language models. They are directly trained on graph datasets with the raw text. The fourth type involves (12) DGTL (Qin et al., 2024), (13) SNS-GPT4 (Li et al., 2024), (14) InstructGraph-INS (Wang et al., 2024), (15) InstructGLM (Ye et al., 2024), (16) GAugLLM (Fang et al., 2024) and (17) Baichuan2 (Yang et al., 2023) six LLM-based methods designed for graph tasks. We report the results from the original papers due to limitations in computational power. For Baichuan2, we directly used it for node class prediction without fine-tuning. Due to the space limitation, we move details on baselines to Appendix A.
Setup. We implement and conduct experiments for LensGNN by PyTorch. We primarily use Baichuan2-13B (Yang et al., 2023) as the backbone LLM for LensGNN. LensGNN ensemble multiple GNN models, including GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), GIN (Xu et al., 2019), each with two layers. We perform a grid search to fine-tune hyperparameters based on the results on the validation set. The specific hyperparameter search space is shown in the Table 2. Some fixed LoRA settings included capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step. We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1. Specifically, we utilize LoRA+ (Hayou et al., 2024). For baselines that reported results on the datasets, we directly report the results from their original papers in the table. Due to computational power limitations, for those datasets where the original papers of the baselines did not report results, we left them blank in the table. We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
Table 4. Ablation study including ‘GNN Encoder’, ‘Alignment’, ‘With Text’ and ‘With Neighbor’.
[…]
These elements enhance the model’s ability to integrate representations and contextual information, resulting in higher accuracy across various datasets. GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%) - - Yes 1 87.82 93.77 74.47 GCN - Yes 1 89.29 94.23 74.35 GAT - Yes 1 90.03 94.26 73.37 GIN - Yes 1 88.92 94.82 73.26 GCN, GAT, GIN Yes No 1 19.92 40.06 34.92 GCN, GAT, GIN No Yes 1 90.77 93.96 73.73 GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03 GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
[…]
Model Bert-base T5-base(Encoder only) T5-base Falcon-7b InternLM2.5-7b-chat Baichuan2-13B
Parameters 110M 110M 220M 7B 7B 13B
Cora With GCN 81.48 78.52 76.10 85.23 89.29 89.29
With GAT 81.85 79.26 77.94 80.44 89.66 90.03
With GIN 87.50 88.24 88.24 87.45 89.29 88.92
Ensemble All 80.37 80.00 77.57 89.66 90.03 90.40
Pubmed With GCN 94.67 94.58 94.03 94.67 94.72 94.23
With GAT 94.37 94.53 94.19 94.37 95.13 94.26
With GIN 94.01 94.17 94.00 94.52 94.87 94.82
Ensemble All 94.17 93.51 93.95 94.97 95.13 95.68
5.2. Overall Performance Comparison (RQ1)
We conduct experiments to compare LensGNN with 17 state-of-the-art competitors on Cora, PubMed and ogbn-arXiv datasets. For Cora and PubMed, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set. For ogbn-arXiv, we split the dataset as suggested in (Hu et al., 2020). We use pre-trained LM to encode node representations based on their original text for all GNN models. For fairness, we use the same pre-trained LM, SentenceBERT (Reimers, 2019), as in LensGNN. Additionally, we compare our backbone LLMs, Baichuan2 (Yang et al., 2023), on the node classification task, where we use Baichuan2 for node label inference directly without fine-tuning. For LensGNN, we report the results of several variants of LensGNN , including a model that integrates GCN, GAT, and GIN simultaneously, as well as variants that ensemble each pairwise combination of these three GNNs.
Table 3 shows the classification accuracy of all the methods. From the tables, we make the following observations: (1) LensGNN outperforms the competitive methods over all the datasets. (2) Under the condition of using the same pre-trained LM to encode node representations, LensGNN outperforms individual GNN models’ classification performance across all datasets. Particularly on the PubMed dataset, LensGNN’s classification accuracy surpasses that of the best GNN methods by more than 5%. This demonstrates the effectiveness of using LLMs to ensemble multiple GNN models in node classification tasks. (3) The LM-based methods encode node representations using raw text to achieve classification. However, these approaches discard the structural properties unique to graph data, resulting in the loss of valuable information and an inability to achieve optimal performance. This indicates the necessity of structural information in graph-related tasks. (4) Some methods attempt to use LLMs to accomplish various graph tasks. For example, InstructGLM describes graph structures in natural language and feeds them into a LLM to achieve node classification. However, LensGNN still achieved better performance across all datasets because LensGNN can leverage the LLM to ensemble the strength of multiple GNNs. (5) Baichuan2 showed poor performance and could not even outperform the language models. This is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Overall, LensGNN harnesses the strengths of multiple GNNs through LLMs and avoids the tedious effort of cherry-picking GNNs across datasets, thereby achieving the best performance across all datasets.
5.3. Ablation Study (RQ2)
Figure 4. Inference Speed and Accuracy of Our Method Based on Different Backbone LLMs.
In this section, we conduct an ablation study on the Cora, PubMed, and ogbn-arXiv datasets to evaluate the impact of each model component. Specifically, we explore various configurations of the ‘GNN Encoder’ ( using different GNNs ), ‘Alignment’ ( whether multiple GNNs are aligned when training GNNs ), ‘With Text’ ( whether node text is included in prompt) and ‘With Neighbor’ ( the number of neighbour hops of the node text when inputting LLMs ). The results are presented in Table 4. From the table, we have the following observations:
Performance of different GNN encoders. We find that ensembling multiple GNNs noticeably outperforms using a single GNN. For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%. This clearly illustrates that by integrating the strengths of various types of GNNs, LensGNN can effectively enhance overall performance.
Impact of Alignment. On the PubMed dataset, the results for aligned GNN and unaligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively. This indicates that aligning multiple GNNs representation space can help LLMs better understand graph tokens. Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
The Importance of Node Text. It can be observed that node text plays a key role in the supervised fine-tuning of LLMs. Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, enabling LLM grasp structural information more deeply, thereby improving classification performance.
Overall, the ablation study confirms that each component of the model contributes significantly to its overall performance.
5.4. Performance with different LLMs (RQ3)
To further demonstrate the effectiveness of LensGNN, we use other LLMs as backbones, including three smaller LMs: BERT-base (Devlin, 2018), T5-base (Raffel et al., 2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B (Almazrouei et al., 2023) and InternLM2.5-7B-chat (Cai et al., 2024). The experiments based on these LLMs include aligning each language model separately with the representation of each type of GNN, as well as aligning them simultaneously with all three types of GNNs (GCN, GAT, GIN) and performing an ensemble. All results are presented in Table 5.
From the table, we make the following observations: (1) Smaller LMs BERT-base, T5-base, and T5-base (Encoder only) can achieve good results when ensemble a single GNN, but the performance drops significantly when ensemble multiple GNN models. This suggests that models with a smaller parameters lack the capability to integrate multiple GNNs. (2) In contrast, LLMs with large parameters that ensemble a single GNN generally perform better than smaller LMs in most cases. (3) Across different large parameter language models, the ensemble of multiple GNN models achieved the best performance on all datasets. Overall, all the observations indicate that LLMs can better achieve the ensemble of GNN models, and the multi-GNNs ensemble can harnesses the strengths of multiple GNNs to achieve the best predictions.
5.5. Model Efficiency Study (RQ4)
In this section, we analyze the computational efficiency of LensGNN during both the training and inference phases.
Model Training Efficiency. We divide the process of LensGNN into two phases: Aligning multi-GNNs and Ensembling multi-GNNs with LLM. In the first phase, the training efficiency primarily depends on the efficiency of the GNN encoders utilized, while also being related to the hyperparameter Graph token $t$. In the second phase, the training efficiency predominantly relies on the fine-tuning efficiency of the LLMs. We employ LoRA technology to reduce the number of parameters requiring fine-tuning. For instance, in the case of Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters. This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment, enabling a training batch size of 4 per device.
Model Inference Efficiency. During inference, all parameters are frozen. Since the model parameters in the first stage are much smaller than the LLM in the second stage, the time cost in the inference phase mainly comes from the second stage’s LLM inference. In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Cora and Pubmed datasets are illustrated in Figure 4. Overall, the inference speed ranges from 2 to 6 samples per second. Although a higher parameter count in the LLM results in reduced inference speed, it concurrently yields superior accuracy. In practical deployments, the choice of backbone LLM can be made by considering the balance between computational power and application requirements.
5.6. Hyperparameter Sensitivity Analysis (RQ5)
The number of Graph Tokens $t$ is an important hyperparameter, representing how many tokens each node is divided into before being input into LLM. We perform experiments on the Cora and Pubmed dataset with $t$ set to values of 1, 2, 4, 8, 16 and 32, and the results are shown in Figure 5.
The results indicate that as the number of Graph Tokens increases, the model’s performance initially rises before gradually declining. The optimal experimental results are achieved when the number of Graph Tokens is set to 8 on both datasets. When the number of tokens is too small, the distinction between nodes becomes insufficient, making it difficult for LLMs to fully learn the semantic information. On the other hand, when there are too many tokens, the representation of Graph Tokens may introduce noise or redundant information, which interferes with the model’s understanding of the graph representations. Therefore, an appropriate number of tokens helps LLMs effectively capture more complicated structural information, enhancing semantic comprehension and leading to better performance.
[…]
6. CONCLUSION
We studied the problem of how to ensemble multi-GNNs in this paper and proposed LensGNN, which ensembles multi-GNNs with LLMs. LensGNN adopts two phases of alignment: the multi-GNN alignment aims to map node representations from different GNNs into the same low-dimensional space, while the GNN-LLM alignment injects graph representation as graph tokens into LLM. After that, we perform supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology. Finally, we conducted extensive experiments to show the effectiveness of our ensembling model LensGNN. In particular, our experimental results showed that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
Appendix A Baselines
To evaluate the effectiveness of LensGNN, we compare it with the SOTA methods. Details of these baselines are summarized as follows.
(1) MLP: Multilayer Perceptron (MLP) is a type of artificial neural network that consists of multiple layers of nodes (neurons) connected by weights and is primarily used for supervised learning tasks, such as classification.
(2) Graph Neural Networks: GCN (Kipf and Welling, 2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data. GAT (Veličković et al., 2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN (Xu et al., 2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. GraphSAGE (Hamilton et al., 2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data. Graphormer (Ying et al., 2021) is built upon the standard Transformer (Vaswani, 2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks. Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data. GT (Dwivedi and Bresson, 2012) presented a simple yet effective approach to generalize transformer networks on arbitrary graphs. It includes an elegant way to fuse node positional features using Laplacian eigenvectors for graph datasets, inspired from the heavy usage of positional encodings in NLP transformer models and recent research on node positional features in GNNs. CoarFormer (Kuang et al., 2024) is a two-view architecture that captures fine-grained local information using a GNN-based module on the original graph and coarse yet long-range information using a Transformer-based module on the coarse graph.
(3) Language Models: BERT (Devlin, 2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT (Reimers, 2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa (He et al., 2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
(4) LLM Based: DGTL (Qin et al., 2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. SNS-GPT4 (Li et al., 2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. InstructGraph-INS (Wang et al., 2024) is an advanced framework that integrates instruction-based learning with graph neural networks (GNNs) to improve performance in various graph-based tasks. The model leverages both explicit instructions and implicit graph structures to enhance representation learning and decision-making processes. InstructGLM (Ye et al., 2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. GAugLLM (Fang et al., 2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. Baichuan2 (Yang et al., 2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment.
Appendix B GNN Encoders
Graph Convolutional Networks (GCNs) GCNs extend the concept of convolutional neural networks (CNNs) to graph data. They operate by aggregating and transforming feature information from a node’s local neighborhood, effectively capturing the graph’s structural information. The propagation rule in GCNs typically involves normalizing the adjacency matrix to ensure proper scaling of node features. GCNs have been widely used in tasks such as node classification, link prediction, and graph classification due to their ability to learn representations that incorporate both node features and graph topology.
Graph Attention Networks (GATs) GATs introduce an attention mechanism to the graph convolutional framework, allowing the model to learn the relative importance of neighboring nodes when aggregating their features. This is achieved by computing attention coefficients for each pair of connected nodes and using these coefficients to perform a weighted aggregation of the neighbor features. The attention mechanism in GATs makes them more flexible and capable of handling graphs with varying degrees of connectivity and node importance. GATs have shown strong performance in various graph-based tasks, providing a way to dynamically adjust the focus on relevant neighbors.
Graph Isomorphism Networks (GINs) GINs are designed to address the limitations of previous GNN models in distinguishing graph structures. They aim to be as powerful as the Weisfeiler-Lehman graph isomorphism test, which is a method for determining whether two graphs are isomorphic. GINs use a simple yet effective aggregation function that sums the features of a node’s neighbors and applies a MLP to the result. This approach ensures that GINs can capture complex graph structures and provide highly discriminative node and graph embeddings. GINs are particularly effective in graph classification tasks, where capturing subtle differences in graph topology is crucial.
Appendix C Detailed hyperparameters
The detailed hyperparameters of the first and second stages of LensGNN  are shown in Tables 6 and 7, respectively.
Table 6. Detailed hyperparameters of aligning multi-GNNs.
Model GCN GAT GIN
Dataset Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv
Hidden size 256 256 256 256 256 256 256 256 256
GNN Representation size 40960 40960 20480 40960 40960 20480 40960 40960 20480
Number of layers 2 2 2 2 2 2 2 2 2
Learning rate of GNN encoder 0.0002 0.0002 0.01 0.0002 0.0002 0.01 0.0001 0.0001 0.004
Initial learning rate of the classifier 0.0002 0.0002 0.0005 0.0002 0.0002 0.0005 0.0001 0.0001 0.0005
Dropout Rate 0.25 0.25 0.2 0.25 0.25 0.2 0.25 0.25 0.2
Weight Decay 0.25 0.25 0 0.25 0.25 0 0.25 0.25 0
Table 7. Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
Model Baichuan2-13B Falcon-7B InternLM2.5-7B-chat
Dataset Cora Pubmed ogb-arxiv Cora Pubmed Cora Pubmed
Lora dropout 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Loraplus Ir ratio 16 16 16 16 16 16 16
Prompt cutoff length 2047 2047 2047 2047 2047 2047 2047
Training batch size per device 4 4 4 4 4 4 4
Gradient accumulation steps 1 1 1 1 1 1 1
Learning rate 0.0001 0.00005 0.0001 0.0001 0.0001 0.0001 0.0001
Early stop epoch 3 2 3 3 2 3 2
Warmup Ratio 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Parameter precision FP16 FP16 FP16 FP16 FP16 FP16 FP16

### v2

@@
 Can Large Language Models Act as Ensembler for Multi-GNNs?
-Abstract.
+Abstract
 Graph Neural Networks (GNNs) have emerged as powerful models for learning from graph-structured data.
-However, GNNs lack the inherent semantic understanding capability of rich textual nodes attributes, limiting their effectiveness in applications.
+However, GNNs lack the inherent semantic understanding capability of rich textual node attributes, limiting their effectiveness in applications.
 On the other hand, we empirically observe that for existing GNN models, no one can consistently outperforms others across diverse datasets.
@@
 Then, through LoRA fine-tuning, it aligns the space between the GNN and the LLM, injecting graph tokens and textual information into LLMs.
-This allows LensGNN to ensemble multiple GNNs and leverage LLM’s strengths, leading to a deeper understanding of both textual semantic information and graph structural information.
-Experimental results show that LensGNN  outperforms existing models.
-This research advances text-attributed graph ensemble learning by providing a robust, superior solution for integrating semantic and structural information.
+This allows LensGNN to ensemble multiple GNNs and take advantage of the strengths of LLM, leading to a deeper understanding of both textual semantic information and graph structural information.
+The experimental results show that LensGNN  outperforms existing models.
+This research advances text-attributed graph ensemble learning by providing a robust and superior solution for integrating semantic and structural information.
 We provide our code and data here: https://anonymous.4open.science/r/EnsemGNN-E267/.
-Large Language Models, Text-attributed Graph, Graph Neural Networks
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
-1. INTRODUCTION
+Can Large Language Models Act as Ensembler for Multi-GNNs?
+Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
+1 INTRODUCTION
 Graphs are structured data that captures the inter-relations between entities in the real world.
-To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced (Kipf and Welling, 2016) and message passing is the main mechanism for neighborhood aggregation (Hamilton et al., 2017a; Veličković et al., 2017).
-GNNs have demonstrated significant success across various applications such as social network analysis (Hamilton et al., 2017a), recommendation systems (Wang et al., 2019), and molecular property prediction (Gilmer et al., 2017).
+To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced Kipf and Welling (2016) and message passing is the main mechanism for neighborhood aggregation Hamilton et al.
+(2017a); Veličković et al.
+(2017).
+GNNs have demonstrated significant success across various applications such as social network analysis Hamilton et al.
+(2017a), recommendation systems Wang et al.
+(2019), and molecular property prediction Gilmer et al.
+(2017).
+Figure 1: Comparison between GNN models on node classification across different datasets.
 Despite the success, two major challenges remain unresolved.
-First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes (Zhao et al., 2019).
+First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes Zhao et al.
+(2019).
 This can lead to the loss of valuable semantic information during the learning process, limiting the effectiveness of GNNs in applications where node features contain meaningful texts, such as citation networks and social media platforms.
 Second, while there have been a series of GNNs proposed, no one has been shown to consistently lead others across datasets of various domains.
-For example, we compare the node classification performance of four representative GNNs, namely, APPNP (Gasteiger et al., 2018), GAT (Veličković et al., 2017), GCN (Kipf and Welling, 2016) and GIN (Xu et al., 2019) on three benchmark datasets (Yan et al., 2023): Cora, Citeseer and PubMed.
+For example, we compare the node classification performance of four representative GNNs, namely, APPNP Gasteiger et al.
+(2018), GAT Veličković et al.
+(2017), GCN Kipf and Welling (2016) and GIN Xu et al.
+(2019) on three benchmark datasets Yan et al.
+(2023): Cora, Citeseer and PubMed.
 For fairness, we set the equal number of hidden representation layers and also the same dimensionality size.
@@
 This also restricts the wide applicability of GNNs.
-[…]
-Comparison of node classification accuracy for various GNN models across different datasets.
-To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
-Figure 1.
-Comparison of node classification accuracy for various GNN models across different datasets.
-To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
-To remedy the incapability of GNNs in semantic understanding, existing works (Qin et al., 2024; Li et al., 2024; Wang et al., 2024; Ye et al., 2024; Fang et al., 2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation.
+To remedy the incapability of GNNs in semantic understanding, existing works Qin et al.
+(2024); Li et al.
+(2024); Wang et al.
+(2024); Ye et al.
+(2024); Fang et al.
+(2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation.
 The inherent advantage of GNNs in utilizing graph structure, fused with the strength of LLMs in semantic understanding, mutually enhances both models and leads to superior performance in downstream tasks.
@@
 Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets.
-The model captures semantic information from node textual attributes using LLMs, while simultaneously enabling LLMs to implicitly evaluate and combine the strengths of multiple GNNs.
 By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets.
@@
   • We propose a novel method LensGNN to ensemble multi-GNNs with LLMs.
-To our best knowledge, we are the first work to attempt the trial.
-LensGNN combines the strengths of both GNNs and LLMs, which can be used to avoid the intricate process of manual GNN selection on various datasets.
-  • We introduce an effective graph prompt tuning paradigm, which enhances LLMs’ capability in comprehending graph topology with two phases of alignment.
-The multi-GNN alignment maps multi-GNNs into the same low-dimensional space while the GNN-LLM alignment injects graph token representations into LLMs.
-  • We conduct extensive experiments to show the effectiveness of LensGNN.
-In particular, we compare LensGNN with 17 state-of-the-art competitors.
-Our experimental results show that LensGNN outperforms baselines in a majority of cases, demonstrating its superior performance.
-2. RELATED WORK
-GNNs (Kipf and Welling, 2016; Xu et al., 2019; Veličković et al., 2017; Hamilton et al., 2017b) have been widely used in learning from graph-structured data.
-Existing GNN studies have been conducted on various types of graphs such as heterophilic graphs (Zhu et al., 2020; Pei et al., 2020; Bi et al., 2024), signed graphs (Derr et al., 2018; Huang et al., 2019) and temporal graphs (Rossi et al., 2020; Longa et al., 2023; Xu et al., 2020).
-Further, to improve the effectiveness of GNNs, there are also works (Nguyen et al., 2023; Rusch et al., 2023; Black et al., 2023) devoted to solving the notorious over-smoothing (Chen et al., 2020) and over-quashing problems (Akansha, 2023).
-In parallel, to scale GNNs to large-scale graphs, some works (Liao et al., 2022) focus on enhancing the model efficiency.
-A comprehensive GNN-related survey can be found at (Wu et al., 2020).
+  • We introduce an effective graph prompt tuning paradigm with both the multi-GNN alignment and the GNN-LLM alignment.
+  • We conduct extensive experiments to show the superiority of LensGNN.
+2 RELATED WORK
+GNNs Kipf and Welling (2016); Xu et al.
+(2019); Veličković et al.
+(2017); Hamilton et al.
+(2017b) have been widely used in learning from graph-structured data.
 Despite their advantages, GNNs struggle to process textual attributes effectively, as traditional GNNs lack the semantic understanding capability that is necessary to handle text-attributed graphs (TAGs).
-In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) (Zoph and Le, 2017).
+In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) Zoph and Le (2017).
 However, it is very computationally expensive, which necessitates new exploration.
 Recently, LLMs have revolutionized tasks involving semantic understanding and language generation.
-Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies (Jin et al., 2023).
-One primary category of approaches takes LLMs as predictors in graph-based tasks, where models like GraphLLM (Chai et al., 2023) utilizes few-shot learning to address graph reasoning challenges.
-ENGINE (Zhu et al., 2024) is designed to efficiently integrate textual and topological information by adding a tunable lightweight GNN side structure alongside each layer of LLMs.Similarly, heuristic reasoning approaches, including Chain-of-Thought (CoT) (Wei et al., 2022) and models like StructGPT (Jiang et al., 2023), have demonstrated improved graph-based inference capabilities.
-Another notable category is to employ LLMs as encoders, as seen in the works of TextGNN (Zhu et al., 2021) and AdsGNN (Li et al., 2021), which focus on optimization strategies for encoding graph data.
-Employing LLMs as encoders involves using large language models to extract textual features, which are then used as initial feature vectors for nodes or edges in a graph.
-These feature vectors are passed into graph neural networks (GNNs) for further processing, allowing the model to incorporate both the textual information from the LLM and the structural information from the graph.
-Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods such as LTRN (Zhang et al., 2021) and GLEM (Zhao et al., 2022), where iterative training helps improve cross-modality learning.
-LLMs and GNNs are aligned through iterative pseudo-label training (prediction alignment) or by using contrastive learning to align their latent representations (latent space alignment).
-Finally, graph-empowered LLMs like GreaseLM (Zhang et al., 2022) and DRAGON (Yasunaga et al., 2022) illustrate the potential of enhancing language models with graph-based knowledge to enrich semantic understanding.
-They integrate the strengths of large language models and graph structures by modifying Transformer architectures to encode text and graph data simultaneously, allowing for a more comprehensive understanding of the interconnected information within graphs.
+Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al.
+(2023).
+One primary category of approaches takes LLMs as predictors in graph-based tasks, like GraphGPT Tang et al.
+(2024), LLaGA Chen et al.
+(2024) GraphLLM Chai et al.
+(2023) and ENGINE Zhu et al.
+(2024).
+Another notable category is to employ LLMs as encoders, as seen in the works of OFA Liu et al.
+(2024), TextGNN Zhu et al.
+(2021) and AdsGNN Li et al.
+(2021), which focus on optimization strategies for encoding graph data.
+Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods, such as GOFA Kong et al.
+(2024), LTRN Zhang et al.
+(2021) and GLEM Zhao et al.
+(2022).
 These diverse approaches highlight the evolving landscape of LLM and GNN integration.
 However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
-3. PRELIMINARIES
+3 PRELIMINARIES
 Text-attributed graph.
@@
 In the message-passing layer, the representation vector $h_{v}^{(t)}$ of node $v$ is updated after the $t$-th iteration using the following formula:
-(1) $h_{v}^{(t)}=\text{UPDATE}\left(\text{AGGREGATE}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$
+$h_{v}^{(t)}=\text{UPDATE}\left(\text{AGG}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$ (1)
 where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$.
-The AGGREGATE function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state.
+The AGG function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state.
 Through this iterative process, GNNs are able to learn increasingly rich node representations, capturing both local and global structural information in the graph.
-The specific implementations of the AGGREGATE and UPDATE functions can vary depending on the particular GNN model.
-4. METHODOLOGY
-To enable the LLM to integrate multiple GNNs, alignment is primarily required in two aspects.
-First, the output representations of all the GNNs need to be aligned into a common vector space.
-Second, there must be alignment between the output representations of the GNNs and the semantic space required by the LLM.
-Based on this idea, we have designed the model in the following two steps.
-The structure of the whole model is shown in Fig.
-2.
-4.1. Aligning multi-GNNs
+The specific implementations of the AGG and UPDATE functions can vary depending on the particular GNN model.
+4 METHODOLOGY
+This section introduces the main steps in LensGNN, which includes aligning multi-GNNs and ensembling multi-GNNs with LLM. The overall model procedure is given in Fig. 2.
+4.1 Aligning multi-GNNs
 Given multiple GNNs, they could generate node representations in different low-dimensional spaces. Therefore, before ensembling GNNs, we need to first align them.
-Figure 2.
-The overall process of LensGNN.
-Step 1 is the workflow of the pre-training phase of LensGNN’s GNNs.
-We employ a language pre-training encoder to extract semantic information from the textual attributes of nodes.
-Concurrently, we train the GNN encoder to capture structural information.
-During this phase, a MLP is utilized as a classifier, and the model is trained through gradient feedback.
-Step 2 is the workflow of the LoRA (Low-Rank Adaptation) supervised fine-tuning stage of LensGNN’s LLM.
-At this stage, we freeze the parameters of the GNNs and infer the hidden representations of the graph nodes.
-We segment GNN representations and convert them into graph tokens that LLM can read.
-The graph tokens, node text tokens, and other prompts are concatenated to form an instruction, which is then input into the LLMs for LoRA supervised fine-tuning.
-[…]
-In the pre-training phase, LensGNN consists of the following architecture: a pre-trained language encoder with frozen parameters, a GNN layer, a linear mapping layer, and an MLP classifier, all of which have parameters that are updated during training.In the fine-tuning phase of LensGNN, the parameters of the pre-trained language encoder and the GNN layers are frozen, and only the Lora parameters of the LLM are updated.
-In addition, the size of Hidden Representation is (1, t * e), where e is the length of the enbedding of each token in LLM, and t is the number of tokens each node representation is converted into.
+Figure 2: The overall process of LensGNN.
 Node feature initialization.
 To utilize the rich semantic information of textual features, we use pre-trained language models to initialize them.
-Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT (Reimers, 2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
-(2) $\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$
+Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT Reimers (2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
+$\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$ (2)
 This process allows texts of any dataset and length to be converted into the same feature space.
@@
 For the $k$-th GNN, we denote the node embedding matrix $H^{[k]}$ generated by the final layer as:
-(3) $H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$
+$H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$ (3)
 where $\tilde{X}$ is the initial node embedding matrix from Eq.
@@
 Reshape node representations.
-The output of each GNN is then connected to a linear layer, which transforms node representations to the dimensionality of hidden embeddings in LLMs.
-This step paves the way for the subsequent GNN and LLM alignment.
+The output of each GNN is then connected to a linear layer, which transforms node representations into the dimensionality of hidden embeddings in LLMs.
+This step paves the way for the subsequent alignment of GNN and LLM.
 Details will be given in the next section.
 For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
-(4) $\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$
+$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$ (4)
 Note that the shape of $\tilde{h}^{[k]}_{i}$ is a one-dimensional vector of length $t\times e$, where $t$ is a hyperparameter indicating the number of graph tokens each node representation is mapped to.
@@
 Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the integration of extensive graph structural information into the LLM in the subsequent steps.
-4.2. Ensembling multi-GNNs with LLM
+4.2 Ensembling multi-GNNs with LLM
 After GNNs are trained, we can get node representations capturing rich graph structural information.
-Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure (Guo et al., 2023).
+Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure Guo et al.
+(2023).
 Therefore, in the second stage, we empower LLMs to comprehend graph representations.
@@
 However, due to the extensive number of parameters in these models, fine-tuning requires significant computational resources, which is challenging for practical applications.
-As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector (Tang et al., 2024), mapping GNN representations into the semantic space.
+As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector Tang et al.
+(2024), mapping GNN representations into the semantic space.
 While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
 […]
-Figure 3.
-A prompt template.
-The prompt is divided into two parts: Instruction and Input.
-The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content.
-The latter provides the specific node text inputs and their multiple GNN representations.
 Prompt design and Multi-GNN ensembling.
@@
 The prompt design follows three key principles:
-  1. (1)
-    Simultaneous input of textual strings from a node and its adjacent neighbors, graph tokens of the node, and task specification.
-Existing studies (Kipf and Welling, 2016; Veličković et al., 2017; Verma et al., 2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction.
+  1. 1.
+    Simultaneous input of textual strings from target node and its adjacent neighbors, graph tokens of the node, and task specification.
+Existing studies Kipf and Welling (2016); Veličković et al.
+(2017); Verma et al.
+(2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction.
 Further, clear task instructions are necessary for accurate prediction.
-  2. (2)
+  2. 2.
     Differentiation between text tokens and graph tokens.
 It directs LLM to correctly distinguish between textual tokens and graph tokens within different segments of the prompt, thus preventing confusion.
-  3. (3)
+  3. 3.
     Guidance for learning from multi-GNNs: It leads LLM to implicitly learn how to effectively combine strengths of different GNNs.
-Based on the above principles, we provide a prompt template used in our experiments, as shown in Figure 3.
-Details on the prompt are given in the caption of the figure.
-In practice, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously.
+Based on the above principles, we provide a prompt template used in our experiments.
+Details on the prompt template are given in Appendix D.
+Further, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously.
 In our experiments, we merge them into one step.
-5. EXPERIMENT
-In this section, we conduct extensive experiments to evaluate the performance of our model across various datasets, aiming to answer the following questions:
-  • RQ1: How does LensGNN perform compared to SOTA baseline models?
-  • RQ2: What is the importance of different components in LensGNN?
-  • RQ3: Can small language models serve as multi-GNN ensembler?
-  • RQ4: How efficient is our method LensGNN ?
-  • RQ5: Is our method LensGNN insensitive to model hyper-parameters?
-Table 1. Datasets statistics.
-Dataset # Nodes # Edges # Classes
-Cora 2,708 5,429 7
-PubMed 19,717 44,338 3
-ogbn-arXiv 169,343 1,166,243 40
-Table 2. Hyperparameter Search Spaces.
-Hyperparameter Search space
-learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
-dropout $[0,0.25]$
-loraplus lr ratio $\{16.0,24.0,32.0\}$
-Graph token $t$$\{1,2,4,8,16\}$
-Table 3.
-The classification performance over the methods on 3 datasets.
+5 EXPERIMENT
+Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods.
+LensGNN utilizes Baichuan2-13B-chat as the backbone LLM.
 We highlight the best score on each dataset in bold and underline the runner-up’s.
-The - means that we did not obtain results from their original paper.
-Model type Model Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
-MLP MLP 66.42 82.40 61.47
-GNN GCN 85.97 83.78 70.81
-GAT 86.71 83.59 70.05
-GIN 85.60 82.03 67.03
-GraphSAGE 84.87 87.01 70.35
-Graphormer 80.41 88.75 72.81
-GT 86.42 88.24 -
-CoarFormer 88.69 89.75 71.66
-Language Model BERT 80.15 93.91 72.78
-SentenceBERT 78.82 92.49 71.42
-DeBERTa 77.79 93.45 72.90
-LLM Based DGTL 81.10 87.10 -
+Model type Model Cora PubMed ogbn-arXiv Citeseer Wiki-CS
+MLP MLP 66.42 82.40 61.47 71.13 68.41
+GNN GCN 85.97 83.78 70.81 77.74 77.15
+GAT 86.71 83.59 70.05 78.21 78.94
+GIN 85.60 82.03 67.03 75.07 68.97
+GraphSAGE 84.87 87.01 70.35 73.61 79.56
+Graphormer 80.41 88.75 72.81 71.28 72.07
+Language Model BERT 80.15 93.91 72.78 73.17 78.33
+SentenceBERT 78.82 92.49 71.42 72.79 77.92
+DeBERTa 77.79 93.45 72.90 73.13 75.11
+LLM Ensembler LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
+LensGNN-[GCN+GIN] 91.88 94.47 75.89 76.64 83.47
+LensGNN-[GAT+GIN] 88.19 95.43 74.31 78.99 82.98
+LensGNN-ALL 90.40 95.68 75.91 79.31 81.78
+5.1 Experimental settings
+Datasets.
+We use eight benchmark datasets, comprising five node classification datasets: Cora, PubMed, ogbn-arXiv, Citeseer and Wiki-CS, and three molecular graph classification datasets Wu et al.
+(2018): BACE, BBBP and ClinTox.
+Details on these datasets are provided in the Appendix A.
+Baselines.
+We compare our method with 19 baselines: MLP, GNN models, LM-based models and LLM-based models.
+Specifically, GNN models include GCN Kipf and Welling (2016), GAT Veličković et al.
+(2017), GIN Xu et al.
+(2019), GraphSAGE Hamilton et al.
+(2017a) and Graphormer Ying et al.
+(2021).
+LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019) and DeBERTa He et al.
+(2020).
+LLM-based models include DGTL Qin et al.
+(2024), SNS-GPT4 Li et al.
+(2024), GAugLLM Fang et al.
+(2024) and Baichuan2-13B Yang et al.
+(2023), which all employ 13B LLMs.
+We also compare LensGNN with other 7B-LLM-based models: GraphGPT Tang et al.
+(2024), LLaGA Chen et al.
+(2024), OFA Liu et al.
+(2024) and GOFA Kong et al.
+(2024).
+Due to the space limitation, details on these baselines are given in Appendix B.
+Setup.
+We use Baichuan2 Yang et al.
+(2023) and InternLM2.5 Cai et al.
+(2024) as the backbone LLM for LensGNN.
+In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers.
+For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input.
+For more details on experiment setup, refer to Appendix C.
+5.2 Node classification results
+Since the large number of baselines report their results on different datasets and settings, we show the classification results in three tables, respectively.
+Table 1 compares LensGNN with MLP, GNN models and small language models.
+Tables 2 summarizes the results against LLM-based models.
+Table 3 evaluates the performance of LensGNN in the 20-shot setting.
+From these tables, we observe:
+(1) LensGNN consistently outperforms all the GNN models and LM-based methods across all the datasets.
+For GNN models, while they initialize node features with SentenceBERT to capture text semantics, LLMs have much stronger capability in text understanding and generation, which explains the advantage of LensGNN.
+For LM-based models, in addition to the weak expressiveness of small language models, they ignore the rich structural information of graphs, degrading their performance.
+(2) LensGNN achieves better performance than LLM-based models, even in the 20-shot setting.
+In particular, we see that Baichuan2-13B performs very poorly.
+We speculate this is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data.
+Different from other LLM-based models, LensGNN integrates the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
+Table 2: Comparison on classification accuracy (%) with LLM-based models. * denotes the best variant of our model and - means that the results are missing from their original papers.
+Model Cora PubMed ogbn-arXiv
+GraphGPT-MIX-7B - 74.16 64.76
+LLaGA-HO-7B(GENERAL) - 94.45 75.01
+InstructGLM-Llama-7B 87.08 93.84 75.70
+DGTL 81.10 87.10 -
 SNS-GPT4 82.50 93.80 74.40
-InstructGraph-INS 89.33 81.09 -
-InstructGLM 90.77 94.62 75.70
 GAugLLM - 83.68 74.15
 Baichuan2-13B 13.65 36.04 4.79
-LensGNN-[GCN+GAT] 88.56 94.11 75.78
-LensGNN-[GCN+GIN] 91.88 94.47 75.89
-LensGNN-[GAT+GIN] 88.19 95.43 74.31
-LensGNN-ALL 90.40 95.68 75.91
-5.1. Experimental Settings
-Datasets. We use three benchmark datasets whose details are given as follows. Statistics of these datasets are summarized in Table 1.
-Cora is a citation network containing 2,708 research papers.
-These papers are categorized into seven classes.
-In this graph, each node represents a scientific paper, and each edge represents a citation.
-Bag-of-words representation is used as the feature vector for each node, which is also assigned a label indicating its research field.
-The task for this dataset is node classification.
-PubMed is a citation network dataset containing 19,717 research papers from the biomedical field.
-These papers are categorized into three classes, and each paper is represented by a bag-of-words feature vector.
-In this graph, each node represents a scientific paper, and each edge represents a citation.
-The task associated with this dataset is node classification.
-ogbn-arXiv is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv.
-In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships.
-These papers are categorized into 40 subject areas.
-We construct node attributes using the title and abstract of the papers.
-The task associated with this dataset is node classification.
-Baselines.
-We compare LensGNN with 17 baselines, which can be categorized into four types.
-The first type involves (1) MLP.
-The second type involves (2) GCN (Kipf and Welling, 2016), (3) GAT (Veličković et al., 2017), (4) GIN (Xu et al., 2019), (5) GraphSAGE (Hamilton et al., 2017a), (6) Graphormer (Ying et al., 2021), (7) GT (Dwivedi and Bresson, 2012) and (8) CoarFormer (Kuang et al., 2024) seven popular GNNs for graph tasks.
-For fairness, we use the node representations obtained from a pre-trained LM as input to the GNN models for learning node classification.
-The third type involves (9) BERT (Devlin, 2018), (10) SentenceBERT (Reimers, 2019) and (11) DeBERTa (He et al., 2020) three language models.
-They are directly trained on graph datasets with the raw text.
-The fourth type involves (12) DGTL (Qin et al., 2024), (13) SNS-GPT4 (Li et al., 2024), (14) InstructGraph-INS (Wang et al., 2024), (15) InstructGLM (Ye et al., 2024), (16) GAugLLM (Fang et al., 2024) and (17) Baichuan2 (Yang et al., 2023) six LLM-based methods designed for graph tasks.
-We report the results from the original papers due to limitations in computational power.
-For Baichuan2, we directly used it for node class prediction without fine-tuning.
-Due to the space limitation, we move details on baselines to Appendix A.
-Setup.
-We implement and conduct experiments for LensGNN by PyTorch.
-We primarily use Baichuan2-13B (Yang et al., 2023) as the backbone LLM for LensGNN.
-LensGNN ensemble multiple GNN models, including GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), GIN (Xu et al., 2019), each with two layers.
-We perform a grid search to fine-tune hyperparameters based on the results on the validation set.
-The specific hyperparameter search space is shown in the Table 2.
-Some fixed LoRA settings included capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step.
-We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1.
-Specifically, we utilize LoRA+ (Hayou et al., 2024).
-For baselines that reported results on the datasets, we directly report the results from their original papers in the table.
-Due to computational power limitations, for those datasets where the original papers of the baselines did not report results, we left them blank in the table.
-We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
-Table 4. Ablation study including ‘GNN Encoder’, ‘Alignment’, ‘With Text’ and ‘With Neighbor’.
-[…]
-These elements enhance the model’s ability to integrate representations and contextual information, resulting in higher accuracy across various datasets.
-GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%) - - Yes 1 87.82 93.77 74.47 GCN - Yes 1 89.29 94.23 74.35 GAT - Yes 1 90.03 94.26 73.37 GIN - Yes 1 88.92 94.82 73.26 GCN, GAT, GIN Yes No 1 19.92 40.06 34.92 GCN, GAT, GIN No Yes 1 90.77 93.96 73.73 GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03 GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
-[…]
+LensGNN*-InternLM2.5-7B 90.03 95.13 74.24
+LensGNN*-Baichuan2-7B 89.85 95.08 75.88
+LensGNN*-Baichuan2-13B 91.88 95.68 75.91
+Table 3: Classification accuracy in 20-shot setting.
+Model Cora PubMed
+GCN 73.5 68.0
+GAT 72.8 68.1
+GIN 75.7 69.3
+OFA 75.34 77.89
+GOFA 77.08 87.33
+LensGNN*-Baichuan2-7B 78.09 89.19
+LensGNN*-InternLM2.5-7B 79.93 87.75
+5.3 Graph classification results
+Table 4: Graph classification results.
+Model BACE BBBP ClinTox
+ACC AUC ACC AUC ACC AUC
+GCN 74.01 84.92 84.14 79.00 90.93 87.23
+GAT 74.67 80.85 76.34 66.12 91.61 86.37
+GIN 68.09 85.66 83.65 79.02 86.91 85.80
+LensGNN-[GCN+GAT] 73.02 72.67 86.09 81.29 98.99 94.00
+LensGNN-[GCN+GIN] 75.32 74.59 88.04 83.27 99.32 96.00
+LensGNN-[GAT+GIN] 76.31 75.99 86.09 81.64 98.99 95.81
+LensGNN-ALL 80.59 80.33 88.53 85.37 98.99 97.63
+We further evaluate the effectiveness of LensGNN on the graph classification task.
+We use classification accuracy (ACC) and AUC scores as the evaluation metrics.
+The results are given in Table  4.
+From the table, LensGNN outperforms GNN baselines in most cases, which shows that combining GNNs and LLMs is inspiring.
+Further, LensGNN-ALL generally performs better than its variants.
+This verifies that leveraging strengths of multiple GNNs is useful.
+Surprisingly, we find that on BACE, LensGNN and its variants achieve smaller AUC scores than GNN models.
+We step into the datasets and observe that BACE is a label-balanced dataset while BBBP and ClinTox are highly imbalanced.
+Note that, for balanced dataset, ACC is more convincing than AUC score, while AUC score is a better indicator than ACC on imbalanced dataset.
+Therefore, the ACC advantage of LensGNN on BACE and the AUC leads on other two datasets have evidently demonstrated the superiority of our method.
+5.4 Ablation study
+We next systematically conduct an ablation study on Cora, PubMed, and ogbn-arXiv datasets to evaluate the importance of major model components.
+Specifically, we explore various configurations of “GNN Encoder” (using different GNNs ), “Alignment” (whether multiple GNNs are aligned when training GNNs ), “With Text” (whether node text is included in prompt) and “With Neighbor” (whether neighborhood information is included or not).
+The results are presented in Table 5.
+Performance of different GNN encoders.
+Ensembling multiple GNNs noticeably outperforms using a single GNN.
+For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%.
+This illustrates that by integrating the strengths of various GNNs, LensGNN can effectively enhance the overall performance.
+Impact of Alignment.
+On the PubMed dataset, the results for unaligned GNN and aligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively.
+This shows that multi-GNN alignment can help LLMs better understand graph tokens.
+Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
+Importance of Node Text.
+It can be observed that node text plays a key role in the supervised fine-tuning of LLMs.
+Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
+Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, thereby improving classification performance.
+Table 5: Ablation study results.
+GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
+- - Yes 1 87.82 93.77 74.47
+GCN - Yes 1 89.29 94.23 74.35
+GAT - Yes 1 90.03 94.26 73.37
+GIN - Yes 1 88.92 94.82 73.26
+GCN, GAT, GIN Yes No 1 19.92 40.06 34.92
+GCN, GAT, GIN No Yes 1 90.77 93.96 73.73
+GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03
+GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
+5.5 Performance with different backbones
+Table 6: The classification accuracy (%) with different LLMs on Cora and PubMed.
 Model Bert-base T5-base (Encoder only) T5-base Falcon-7B InternLM2.5-7B-chat Baichuan2-13B
@@
 Ensemble All 94.17 93.51 93.95 94.97 95.13 95.68
-5.2. Overall Performance Comparison (RQ1)
-We conduct experiments to compare LensGNN with 17 state-of-the-art competitors on Cora, PubMed and ogbn-arXiv datasets.
-For Cora and PubMed, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set.
-For ogbn-arXiv, we split the dataset as suggested in (Hu et al., 2020).
-We use pre-trained LM to encode node representations based on their original text for all GNN models.
-For fairness, we use the same pre-trained LM, SentenceBERT (Reimers, 2019), as in LensGNN.
-Additionally, we compare our backbone LLMs, Baichuan2 (Yang et al., 2023), on the node classification task, where we use Baichuan2 for node label inference directly without fine-tuning.
-For LensGNN, we report the results of several variants of LensGNN , including a model that integrates GCN, GAT, and GIN simultaneously, as well as variants that ensemble each pairwise combination of these three GNNs.
-Table 3 shows the classification accuracy of all the methods.
-From the tables, we make the following observations: (1) LensGNN outperforms the competitive methods over all the datasets.
-(2) Under the condition of using the same pre-trained LM to encode node representations, LensGNN outperforms individual GNN models’ classification performance across all datasets.
-Particularly on the PubMed dataset, LensGNN’s classification accuracy surpasses that of the best GNN methods by more than 5%.
-This demonstrates the effectiveness of using LLMs to ensemble multiple GNN models in node classification tasks.
-(3) The LM-based methods encode node representations using raw text to achieve classification.
-However, these approaches discard the structural properties unique to graph data, resulting in the loss of valuable information and an inability to achieve optimal performance.
-This indicates the necessity of structural information in graph-related tasks.
-(4) Some methods attempt to use LLMs to accomplish various graph tasks.
-For example, InstructGLM describes graph structures in natural language and feeds them into a LLM to achieve node classification.
-However, LensGNN still achieved better performance across all datasets because LensGNN can leverage the LLM to ensemble the strength of multiple GNNs.
-(5) Baichuan2 showed poor performance and could not even outperform the language models.
-This is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data.
-Overall, LensGNN harnesses the strengths of multiple GNNs through LLMs and avoids the tedious effort of cherry-picking GNNs across datasets, thereby achieving the best performance across all datasets.
-5.3. Ablation Study (RQ2)
-Figure 4. Inference Speed and Accuracy of Our Method Based on Different Backbone LLMs.
-In this section, we conduct an ablation study on the Cora, PubMed, and ogbn-arXiv datasets to evaluate the impact of each model component.
-Specifically, we explore various configurations of the ‘GNN Encoder’ ( using different GNNs ), ‘Alignment’ ( whether multiple GNNs are aligned when training GNNs ), ‘With Text’ ( whether node text is included in prompt) and ‘With Neighbor’ ( the number of neighbour hops of the node text when inputting LLMs ).
-The results are presented in Table 4.
-From the table, we have the following observations:
-Performance of different GNN encoders.
-We find that ensembling multiple GNNs noticeably outperforms using a single GNN.
-For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%.
-This clearly illustrates that by integrating the strengths of various types of GNNs, LensGNN can effectively enhance overall performance.
-Impact of Alignment.
-On the PubMed dataset, the results for aligned GNN and unaligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively.
-This indicates that aligning multiple GNNs representation space can help LLMs better understand graph tokens.
-Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
-The Importance of Node Text.
-It can be observed that node text plays a key role in the supervised fine-tuning of LLMs.
-Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
-Role of Neighbors.
-The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, enabling LLM grasp structural information more deeply, thereby improving classification performance.
-Overall, the ablation study confirms that each component of the model contributes significantly to its overall performance.
-5.4. Performance with different LLMs (RQ3)
-To further demonstrate the effectiveness of LensGNN, we use other LLMs as backbones, including three smaller LMs: BERT-base (Devlin, 2018), T5-base (Raffel et al., 2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B (Almazrouei et al., 2023) and InternLM2.5-7B-chat (Cai et al., 2024).
-The experiments based on these LLMs include aligning each language model separately with the representation of each type of GNN, as well as aligning them simultaneously with all three types of GNNs (GCN, GAT, GIN) and performing an ensemble.
-All results are presented in Table 5.
-From the table, we make the following observations: (1) Smaller LMs BERT-base, T5-base, and T5-base (Encoder only) can achieve good results when ensemble a single GNN, but the performance drops significantly when ensemble multiple GNN models.
-This suggests that models with a smaller parameters lack the capability to integrate multiple GNNs.
-(2) In contrast, LLMs with large parameters that ensemble a single GNN generally perform better than smaller LMs in most cases.
-(3) Across different large parameter language models, the ensemble of multiple GNN models achieved the best performance on all datasets.
-Overall, all the observations indicate that LLMs can better achieve the ensemble of GNN models, and the multi-GNNs ensemble can harnesses the strengths of multiple GNNs to achieve the best predictions.
-5.5. Model Efficiency Study (RQ4)
-In this section, we analyze the computational efficiency of LensGNN during both the training and inference phases.
-Model Training Efficiency.
-We divide the process of LensGNN into two phases: Aligning multi-GNNs and Ensembling multi-GNNs with LLM.
-In the first phase, the training efficiency primarily depends on the efficiency of the GNN encoders utilized, while also being related to the hyperparameter Graph token $t$.
-In the second phase, the training efficiency predominantly relies on the fine-tuning efficiency of the LLMs.
-We employ LoRA technology to reduce the number of parameters requiring fine-tuning.
-For instance, in the case of Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters.
-This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment, enabling a training batch size of 4 per device.
-Model Inference Efficiency.
-During inference, all parameters are frozen.
-Since the model parameters in the first stage are much smaller than the LLM in the second stage, the time cost in the inference phase mainly comes from the second stage’s LLM inference.
-In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Cora and Pubmed datasets are illustrated in Figure 4.
-Overall, the inference speed ranges from 2 to 6 samples per second.
-Although a higher parameter count in the LLM results in reduced inference speed, it concurrently yields superior accuracy.
-In practical deployments, the choice of backbone LLM can be made by considering the balance between computational power and application requirements.
-5.6. Hyperparameter Sensitivity Analysis (RQ5)
-The number of Graph Tokens $t$ is an important hyperparameter, representing how many tokens each node is divided into before being input into LLM.
-We perform experiments on the Cora and Pubmed dataset with $t$ set to values of 1, 2, 4, 8, 16 and 32, and the results are shown in Figure 5.
-The results indicate that as the number of Graph Tokens increases, the model’s performance initially rises before gradually declining.
-The optimal experimental results are achieved when the number of Graph Tokens is set to 8 on both datasets.
-When the number of tokens is too small, the distinction between nodes becomes insufficient, making it difficult for LLMs to fully learn the semantic information.
-On the other hand, when there are too many tokens, the representation of Graph Tokens may introduce noise or redundant information, which interferes with the model’s understanding of the graph representations.
-Therefore, an appropriate number of tokens helps LLMs effectively capture more complicated structural information, enhancing semantic comprehension and leading to better performance.
-[…]
-6. CONCLUSION
+To study the impact of backbone models on LensGNN, we use different backbones, including three small LMs: BERT-base Devlin (2018), T5-base Raffel et al.
+(2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B Almazrouei et al.
+(2023) and InternLM2.5-7B-chat Cai et al.
+(2024).
+All the results are presented in Table 6.
+From the table, we see that
+(1) Small LMs BERT-base, T5-base, and T5-base (Encoder only) perform well when applied to a single GNN, but the performance drops significantly when ensembling multiple GNN models.
+This suggests that models with less parameters lack the capability to integrate multiple GNNs.
+(2) For LLMs, they generally perform better than LM backbones.
+For example, for GAT on Cora, the best result for LM backbones is 81.85%, while that for LLM backbones is 90.03%.
+This is because LLMs can provide richer text semantics.
+Further, compared with single GNN, the ensemble of multiple GNNs with LLMs leads to better performance, which verifies that unifying the strengths of multiple GNNs is beneficial for node classification.
+5.6 Model efficiency study
+Training Efficiency.
+The training of LensGNN includes two main phases: aligning multi-GNNs and ensembling multi-GNNs with LLM.
+In the first phase, the time complexity depends on the GNN encoders.
+Since the used GNNs have a linear time complexity w.r.t.
+the number of nodes in the graph, the training process is efficient.
+In the second phase, fine-tuning LLMs is the main computational cost.
+We employ LoRA to reduce the number of parameters to be fine-tuned.
+For example, for Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters.
+This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment.
+Compared to GraphGPT Tang et al.
+(2024), the increase of our training computational cost primarily stems from employing more than one GNN.
+However, in contrast to LLM, the number of parameters in GNN is considerably smaller, which increases marginal training cost.
+Inference Efficiency.
+During inference, all the parameters are frozen.
+The major cost of model inference comes from LLMs.
+In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Pubmed dataset is illustrated in Figure 3.
+Overall, the inference speed ranges from $2$ to $4$ samples per second.
+Although a larger LLM results in reduced inference speed, it concurrently yields superior accuracy.
+Figure 3: Inference speed and accuracy comparison.
+5.7 Hyperparameter sensitivity analysis
+We end this section with model sensitivity analysis on the number of graph tokens $t$, which represents how many tokens each node is mapped into before fed into LLM.
+We conduct experiments on Cora and Pubmed with varying $t$ values, and the results are shown in Figure 4.
+From the figure, we see that, as the number of graph tokens increases, the model’s performance first rises and then decreases.
+For both datasets, the best results are achieved when $t=8$.
+When $t$ is small, graph token representations cannot well capture the semantic information.
+On the other hand, when $t$ is large, the representation of graph tokens could be noise-corrupted, which adversely affects the model’s understanding on graph structure.
+Therefore, an appropriate number of tokens helps LLMs effectively capture structural information, enhancing semantic comprehension and leading to better performance.
+Figure 4: Hyperparameter sensitivity analysis.
+6 CONCLUSION
 We studied the problem of how to ensemble multi-GNNs in this paper and proposed LensGNN, which ensembles multi-GNNs with LLMs.
@@
 In particular, our experimental results showed that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
-Appendix A Baselines
+7 LIMITATIONS
+Our method could suffer from several limitations:
+First, due to limited computational resources, we do not conduct experiments on LLMs with parameters exceeding 13B.
+This implies that the upper limit of our method’s capabilities has not been fully investigated.
+For larger-scale LLMs that are utilized in production environments, they could further improve our model’s performance.
+Second, our method relies on manually crafted prompts to ensemble multi-GNNs. Future studies on prompt design could be a promising direction.
+Third, we only evaluated the performance based on node classification and graph classification tasks.
+Evaluation on other tasks, such as graph dataset generation and graph task interpretability analysis, are valuable research questions.
+8 ETHICAL STATEMENT
+Our work falls under basic research and is not tied to specific applications; therefore, whether our method will be abused and cause negative social impacts depends on the specific applications in which others use our method.
+In addition, our work does not involve any stakeholders benefiting or being disadvantaged, nor does it involve vulnerable groups.
+The datasets we used are all commonly used public datasets, posing no privacy risks, and aligned with their intention for scientific research.
+For the datasets, pre-trained models, and training frameworks utilized, we adhered to their respective open-source licenses: community license for Baichuan-2 Yang et al.
+(2023), Apache License 2.0 for InternLM2.5 Cai et al.
+(2024) , Falcon-7B Almazrouei et al.
+(2023) and LLaMA-Factory Zheng et al.
+(2024)
+Appendix A Datasets
+We use eight benchmark datasets, comprising five node classification datasets: Cora, PubMed, ogbn-arXiv, Citeseer and Wiki-CS, and three molecular graph classification datasets: BACE, BBBP and ClinTox.
+Statistics of these datasets are summarized in Table 7.
+Cora McCallum et al.
+(2000) is a standard citation network dataset consisting of 2,708 research papers in the field of machine learning.
+These papers are categorized into seven classes.
+The title and abstract of each paper are utilized as the textual attributes of the nodes.
+The dataset includes 5,429 citation links, constructing a graph structure among the papers.
+The Cora dataset is commonly used for node classification and evaluating the performance of graph neural network models.
+Citeseer Giles et al.
+(1998) is a citation network dataset comprising 3,312 research papers primarily from the fields of computer science and information technology.
+These papers are categorized into six classes based on their research areas.
+The title and abstract of each paper are used as the textual attributes of the nodes, providing a semantic representation of the content.
+The dataset includes 4,732 citation links, forming a graph structure that represents the citation relationships between the papers.
+The Citeseer dataset is widely used for node classification tasks, and it serves as a benchmark for evaluating the performance of graph neural network models.
+PubMed Sen et al.
+(2008) is another citation network dataset containing 19,717 research papers from the biomedical field.
+These papers are categorized into three classes.
+The title and abstract of each paper are utilized as the textual attributes of the nodes.
+The dataset includes 44,338 citation links, constructing a graph structure among the papers.
+The PubMed dataset is also used for node classification tasks, particularly in testing large-scale graph neural network models.
+ogbn-arXiv Hu et al.
+(2020) is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv.
+These papers are time-ordered by their submission dates and categorized into 40 subject areas.
+In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships.
+The dataset includes 1,166,243 citation links, constructing a citation network among the papers.
+The ogbn-arXiv dataset is commonly used for node classification and studying time-sensitive graph neural network models.
+Wiki-CS Mernyei and Cangea (2022) is derived from Wikipedia and consists of 11,701 web pages related to computer science topics.
+These pages are categorized into 10 classes, each representing a different area of computer science such as artificial intelligence, computer architecture, and software engineering.
+The text content of each page is used as the textual attributes of the nodes, capturing the thematic essence of the pages.
+The dataset includes 216,123 hyperlinks, which construct a graph structure connecting the web pages.
+The Wiki-CS dataset is often utilized for node classification tasks and for evaluating graph-based machine learning models.
+BACE Wu et al.
+(2018) is a collection of inhibitors of human beta-secretase 1 (BACE-1) and provides both quantitative IC50 values and qualitative binary labels indicating the binding results.
+The BACE dataset comprises 1,513 molecular graphs for the molecular property prediction.
+BBBP Wu et al.
+(2018) is used for predicting blood-brain barrier permeability (BBBP), which is crucial for determining whether a molecule can cross the blood-brain barrier.
+The BBBP dataset contains 2,039 molecular graphs for the binary graph classification task.
+ClinTox Wu et al.
+(2018) is a collection of drugs approved by the FDA and those that have failed clinical trials due to toxicity reasons.
+The dataset encompasses two classification tasks for 1,491 drug molecules: (1) clinical trial toxicity and (2) FDA approval status.
+For Cora, PubMed, Citeseer, BACE, BBBP and ClinTox, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set.
+For ogbn-arXiv we split the dataset as suggested in Hu et al. (2020).
+For Wiki-CS, we split the dataset as suggested in Mernyei and Cangea (2022).
+Table 7: Datasets statistics.
+Dataset #Graphs Avg.# Nodes Avg.# Edges # Classes
+Cora 1 2,708 5,429 7
+PubMed 1 19,717 44,338 3
+ogbn-arXiv 1 169,343 1,166,243 40
+Citeseer 1 3,312 4,732 6
+Wiki-CS 1 11,701 216,123 10
+BACE 1,513 34.1 73.7 1
+BBBP 2,039 23.9 51.6 1
+ClinTox 1,491 26.1 55.5 2
+Appendix B Baselines
 To evaluate the effectiveness of LensGNN, we compare it with the SOTA methods. Details of these baselines are summarized as follows.
 (1) MLP: Multilayer Perceptron (MLP) is a type of artificial neural network that consists of multiple layers of nodes (neurons) connected by weights and is primarily used for supervised learning tasks, such as classification.
-(2) Graph Neural Networks: GCN (Kipf and Welling, 2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data.
-GAT (Veličković et al., 2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them.
-GIN (Xu et al., 2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test.
-GraphSAGE (Hamilton et al., 2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data.
-Graphormer (Ying et al., 2021) is built upon the standard Transformer (Vaswani, 2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks.
+(2) Graph Neural Networks: GCN Kipf and Welling (2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data.
+GAT Veličković et al.
+(2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them.
+GIN Xu et al.
+(2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test.
+GraphSAGE Hamilton et al.
+(2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data.
+Graphormer Ying et al.
+(2021) is built upon the standard Transformer Vaswani (2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks.
 Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data.
-GT (Dwivedi and Bresson, 2012) presented a simple yet effective approach to generalize transformer networks on arbitrary graphs.
-It includes an elegant way to fuse node positional features using Laplacian eigenvectors for graph datasets, inspired from the heavy usage of positional encodings in NLP transformer models and recent research on node positional features in GNNs.
-CoarFormer (Kuang et al., 2024) is a two-view architecture that captures fine-grained local information using a GNN-based module on the original graph and coarse yet long-range information using a Transformer-based module on the coarse graph.
-(3) Language Models: BERT (Devlin, 2018) is a groundbreaking model in natural language processing.
+(3) LM Based Models: BERT Devlin (2018) is a groundbreaking model in natural language processing.
 It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously.
 This bidirectional approach enables BERT to capture nuanced meanings and relationships in text.
-SentenceBERT (Reimers, 2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings.
+SentenceBERT Reimers (2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings.
 It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering.
-DeBERTa (He et al., 2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
-(4) LLM Based: DGTL (Qin et al., 2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs.
-SNS-GPT4 (Li et al., 2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications.
+DeBERTa He et al.
+(2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
+Figure 5: A prompt template for node classification task.
+Figure 6: A prompt template for graph classification task.
+(4) LLM Based Models: GraphGPT Tang et al.
+(2024) is a novel framework that integrates Large Language Models (LLMs) with graph structural knowledge through graph instruction tuning.
+This innovative approach enables LLMs to comprehend and interpret the structural components of graphs, thereby demonstrating superior generalization in both supervised and zero-shot graph learning tasks.
+LLaGA Chen et al.
+(2024) integrates the capabilities of LLMs with graph-structured data, enabling LLMs to handle complex graph tasks effectively.
+LLaGA achieves this by reorganizing graph nodes into structure-aware sequences and mapping them into the token embedding space through a versatile projector, demonstrating superior versatility, generalizability, and interpretability across various datasets and tasks.
+InstructGLM Ye et al.
+(2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning.
+DGTL Qin et al.
+(2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs.
+SNS-GPT4 Li et al.
+(2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications.
 It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily.
-InstructGraph-INS (Wang et al., 2024) is an advanced framework that integrates instruction-based learning with graph neural networks (GNNs) to improve performance in various graph-based tasks.
-The model leverages both explicit instructions and implicit graph structures to enhance representation learning and decision-making processes.
-InstructGLM (Ye et al., 2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning.
-GAugLLM (Fang et al., 2024) is a novel framework for augmenting TAGs.
+GAugLLM Fang et al.
+(2024) is a novel framework for augmenting TAGs.
 It leverages advanced large language models like Mistral to enhance self-supervised graph learning.
-Baichuan2 (Yang et al., 2023) refers to the second iteration of the Baichuan large language model (LLM).
+Baichuan2 Yang et al.
+(2023) refers to the second iteration of the Baichuan large language model (LLM).
 It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances.
 Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment.
-Appendix B GNN Encoders
-Graph Convolutional Networks (GCNs) GCNs extend the concept of convolutional neural networks (CNNs) to graph data.
-They operate by aggregating and transforming feature information from a node’s local neighborhood, effectively capturing the graph’s structural information.
-The propagation rule in GCNs typically involves normalizing the adjacency matrix to ensure proper scaling of node features.
-GCNs have been widely used in tasks such as node classification, link prediction, and graph classification due to their ability to learn representations that incorporate both node features and graph topology.
-Graph Attention Networks (GATs) GATs introduce an attention mechanism to the graph convolutional framework, allowing the model to learn the relative importance of neighboring nodes when aggregating their features.
-This is achieved by computing attention coefficients for each pair of connected nodes and using these coefficients to perform a weighted aggregation of the neighbor features.
-The attention mechanism in GATs makes them more flexible and capable of handling graphs with varying degrees of connectivity and node importance.
-GATs have shown strong performance in various graph-based tasks, providing a way to dynamically adjust the focus on relevant neighbors.
-Graph Isomorphism Networks (GINs) GINs are designed to address the limitations of previous GNN models in distinguishing graph structures.
-They aim to be as powerful as the Weisfeiler-Lehman graph isomorphism test, which is a method for determining whether two graphs are isomorphic.
-GINs use a simple yet effective aggregation function that sums the features of a node’s neighbors and applies a MLP to the result.
-This approach ensures that GINs can capture complex graph structures and provide highly discriminative node and graph embeddings.
-GINs are particularly effective in graph classification tasks, where capturing subtle differences in graph topology is crucial.
-Appendix C Detailed hyperparameters
-The detailed hyperparameters of the first and second stages of LensGNN  are shown in Tables 6 and 7, respectively.
-Table 6. Detailed hyperparameters of aligning multi-GNNs.
-Model GCN GAT GIN
-Dataset Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv
-Hidden size 256 256 256 256 256 256 256 256 256
-GNN Representation size 40960 40960 20480 40960 40960 20480 40960 40960 20480
-Number of layers 2 2 2 2 2 2 2 2 2
-Learning rate of GNN encoder 0.0002 0.0002 0.01 0.0002 0.0002 0.01 0.0001 0.0001 0.004
-Initial learning rate of the classifier 0.0002 0.0002 0.0005 0.0002 0.0002 0.0005 0.0001 0.0001 0.0005
-Dropout Rate 0.25 0.25 0.2 0.25 0.25 0.2 0.25 0.25 0.2
-Weight Decay 0.25 0.25 0 0.25 0.25 0 0.25 0.25 0
-Table 7. Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
-Model Baichuan2-13B Falcon-7B InternLM2.5-7B-chat
-Dataset Cora Pubmed ogb-arxiv Cora Pubmed Cora Pubmed
-Lora dropout 0.1 0.1 0.1 0.1 0.1 0.1 0.1
-Loraplus Ir ratio 16 16 16 16 16 16 16
-Prompt cutoff length 2047 2047 2047 2047 2047 2047 2047
-Training batch size per device 4 4 4 4 4 4 4
-Gradient accumulation steps 1 1 1 1 1 1 1
-Learning rate 0.0001 0.00005 0.0001 0.0001 0.0001 0.0001 0.0001
-Early stop epoch 3 2 3 3 2 3 2
-Warmup Ratio 0.1 0.1 0.1 0.1 0.1 0.1 0.1
-Parameter precision FP16 FP16 FP16 FP16 FP16 FP16 FP16
+OFA Liu et al.
+(2024) is a pioneering framework that unifies various graph tasks into a single model, enabling it to address classification tasks across different domains and tasks.
+OFA achieves this by converting graph data into text-attributed graphs (TAGs), using language models to encode diverse text attributes into a common embedding space, and introducing a novel graph prompting paradigm for in-context learning without fine-tuning.
+GOFA Kong et al.
+(2024) is a novel graph foundation model that integrates the strengths of large language models (LLMs) and graph neural networks (GNNs) to enable joint graph and language modeling.
+It achieves this by interleaving GNN layers into a frozen pre-trained LLM, allowing for self-supervised pretraining on graph data and fluidity in handling various graph-related tasks.
+Table 8: Hyperparameter Search Spaces.
+Hyperparameter Search space
+learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
+dropout $[0,0.25]$
+loraplus lr ratio $\{16.0,24.0,32.0\}$
+Graph token $t$$\{1,2,4,8,16\}$
+Table 9: Detailed hyperparameters of aligning multi-GNNs.
+Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
+GCN Cora 256 40960 2 0.0002 0.0002 0.25 0.001
+Pubmed 256 40960 2 0.0002 0.0002 0.25 0.001
+ogbn-arxiv 256 20480 2 0.01 0.0005 0.2 0
+Citeseer 256 40960 2 0.0001 0.0005 0.25 0.001
+Wiki-CS 256 40960 2 0.0002 0.0005 0.25 0.001
+BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
+BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
+ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
+GAT Cora 256 40960 2 0.0002 0.0002 0.25 0.001
+Pubmed 256 40960 2 0.0002 0.0002 0.25 0.001
+ogbn-arxiv 256 20480 2 0.01 0.0005 0.2 0
+Citeseer 256 40960 2 0.0001 0.0005 0.25 0.001
+Wiki-CS 256 40960 2 0.0002 0.0005 0.25 0.001
+BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
+BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
+ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
+GIN Cora 256 40960 2 0.0001 0.0001 0.25 0.001
+Pubmed 256 40960 2 0.0001 0.0001 0.25 0.001
+ogbn-arxiv 256 20480 2 0.004 0.0005 0.2 0
+Citeseer 256 40960 2 0.00005 0.0005 0.25 0.001
+Wiki-CS 256 40960 2 0.0001 0.0005 0.25 0.001
+BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
+BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
+ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
+Table 10: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
+Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
+Baichuan2-13B Cora 0.1 16 4 0.0001 3 0.1
+Pubmed 0.1 16 4 0.00005 2 0.1
+ogbn-arxiv 0.1 16 4 0.0001 3 0.1
+Citeseer 0.15 16 4 0.0001 5 0.1
+Wiki-CS 0.1 16 4 0.0001 5 0.1
+BACE 0.1 16 4 0.0001 4 0.1
+BBBP 0.1 16 4 0.0001 3 0.1
+ClinTox 0.1 16 4 0.0001 3 0.1
+Baichuan2-7B Cora 0.15 16 4 0.0001 3 0.1
+Pubmed 0.15 16 4 0.0001 2 0.1
+ogbn-arxiv 0.15 16 4 0.0001 3 0.1
+InternLM2.5-7B Cora 0.1 16 4 0.0001 3 0.1
+Pubmed 0.1 16 4 0.0001 2 0.1
+ogbn-arxiv 0.1 16 4 0.0001 3 0.1
+Falcon-7B Cora 0.1 16 4 0.0001 3 0.1
+Pubmed 0.1 16 4 0.0001 2 0.1
+Appendix C Experiment setup
+We implement LensGNN by PyTorch.
+We primarily use Baichuan2 Yang et al.
+(2023) and InternLM2.5 Cai et al.
+(2024) as the backbone LLM for LensGNN.
+In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers.
+We perform grid search to fine-tune hyperparameters based on the validation set.
+Details on the search space is given in Table 8.
+We utilize LoRA+ Hayou et al.
+(2024) for fine-tuning.
+Some fixed LoRA settings include capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step.
+We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1.
+For the training of LLMs, we utilize LLaMA-Factory Zheng et al.
+(2024) as a framework.
+For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input.
+For baselines that report results on the adopted datasets, we directly report the results from their original papers.
+For those whose results are missing, we leave them blank.
+For Baichuan2, we directly use it without fine-tuning.
+We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
+Appendix D Prompt design
+The prompt is divided into two parts: Instruction and Input.
+The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content.
+The latter provides the specific text of target node or graph and their multiple GNN representations.
+An example of a prompt template can be seen in Figure 5.
+For graph classification, we follow MolXPT Liu et al.
+(2023) to design prompts for molecular graphs as shown in Figure 6.
+Appendix E Detailed hyperparameter settings
+The details on the setting of hyperparameters in the first and second stages of LensGNN  are shown in Tables 9 and 10, respectively.
+The hyperparameters for the 20-shot experiment are shown in Tables 11 and 12.
+Table 11: Detailed hyperparameters of aligning multi-GNNs (20-shot).
+Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
+GCN Cora 256 32768 2 0.0002 0.0005 0.25 0.001
+Pubmed 256 32768 2 0.0002 0.0005 0.25 0.001
+GAT Cora 256 32768 2 0.0002 0.0005 0.25 0.001
+Pubmed 256 32768 2 0.0002 0.0005 0.25 0.001
+GIN Cora 256 32768 2 0.0001 0.0005 0.25 0.001
+Pubmed 256 32768 2 0.0001 0.0005 0.25 0.001
+Table 12: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune) (20-shot).
+Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
+Baichuan2-7B Cora 0.15 16 4 0.0001 8 0.1
+Pubmed 0.15 16 4 0.0001 9 0.1
+InternLM2.5-7B Cora 0.1 16 4 0.0001 8 0.1
+Pubmed 0.05 16 4 0.0001 9 0.1

### v3

@@
 Can Large Language Models Act as Ensembler for Multi-GNNs?
-Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
+Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Yao Liu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
 1 INTRODUCTION
@@
 In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model.
-Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets.
+Given multiple GNNs, the model ensembles GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets.
 By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets.
@@
 Recently, LLMs have revolutionized tasks involving semantic understanding and language generation.
-Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al.
+Meanwhile, the ensemble of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al.
 (2023).
@@
 (2022).
-These diverse approaches highlight the evolving landscape of LLM and GNN integration.
+These diverse approaches highlight the evolving landscape of LLM and GNN ensemble.
 However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
@@
 For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
-$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$ (4)
+$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\\$ (4)
 Note that the shape of $\tilde{h}^{[k]}_{i}$ is a one-dimensional vector of length $t\times e$, where $t$ is a hyperparameter indicating the number of graph tokens each node representation is mapped to.
@@
 Multi-GNN alignment.
-To align node representations from multiple GNNs, we next feed them into a shared linear layer, which serves as the classifier.
+To align node representations from multiple GNNs, we next feed them into a shared MLP, which serves as the classifier.
 During the training time, we alternatively input representations from different GNNs into the classifier and train them sequentially.
-This training approach allows node representations from multi-GNNs to be aligned, which facilitates the integration of multi-GNNs.
+This training approach allows node representations from multi-GNNs to be aligned, which facilitates the ensemble of multi-GNNs.
+For better alignment, we freeze this MLP starting from a certain training epoch.
 After training, the parameters of GNNs and the linear layer in Eq.
@@
 By incorporating the node feature initialization step that utilizes language pre-trained models, our model gains the ability to extract semantic and structural information from the nodes in the text graph.
-Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the integration of extensive graph structural information into the LLM in the subsequent steps.
+Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the ensemble of extensive graph structural information into the LLM in the subsequent steps.
 4.2 Ensembling multi-GNNs with LLM
@@
 While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
-[…]
+To address the problem, recent advancements in LoRA (Low-Rank Adaptation) Hu et al.
+(2021) training have been proposed to directly fine-tune LLMs for alignment.
+LoRA is an efficient fine-tuning technique for large pre-trained models, and introduces low-rank matrices to simulate the effects of full parameter tuning, significantly reducing the number of training parameters and computational resources required.
+This allows for effective customization of LLMs even with limited resources.
+Thus, we adopt LoRA training to fine-tune the LLMs, enabling them to comprehend GNN representations.
+Specifically, given an instruction prompt, we reserve several dummy tokens in the prompt text at the appropriate positions where the GNN representations need to be inserted.
+For example, in the prompt ‘(GNN type: GCN, representations: $\text{graph\_token}_{1}>,<\text{graph\_token}_{2}>,\ldots,<\text{graph\_token}_{t}>$ ), which category does it belong to?’, $\text{graph\_token}_{i}>$ denotes the $i$-th dummy token.
+Then, for each token in the prompt text, LLM can generate an embedding vector of size $e$ .
+These vectors can form an embedding for the whole prompt text in the format of $\ldots,E^{\prime}_{1},E^{\prime}_{2},\ldots,E^{\prime}_{t},\ldots]$, where $E^{\prime}_{i}$ is the embedding vector of $\text{graph\_token>}_{i}$.
+After that, we overwrite the embeddings of dummy tokens by GNN representations.
+From Eq.
+4, we can generate a GNN representation vector of length $t\times e$, which can be further segmented into $t$ one-dimensional vectors of length $e$, denoted as $E^{G}_{1},E^{G}_{2},\ldots,E^{G}_{t}]$.
+These vectors serve as the embeddings of $t$ graph tokens.
+After substitution, the raw embedding vector of the prompt text becomes $\ldots,E^{G}_{1},E^{G}_{2},\ldots,E^{G}_{t},\ldots]$.
+We next feed it into LLM and use LoRA for fine-tuning.
+In this way, we can automatically leverage the language understanding capability of LLM, and implicitly align structure and semantics.
 Prompt design and Multi-GNN ensembling.
@@
 Our goal is to leverage the strengths of both multi-GNNs and LLM to output more accurate predictions.
-The prompt design follows three key principles:
-  1. 1.
-    Simultaneous input of textual strings from target node and its adjacent neighbors, graph tokens of the node, and task specification.
-Existing studies Kipf and Welling (2016); Veličković et al.
-(2017); Verma et al.
-(2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction.
-Further, clear task instructions are necessary for accurate prediction.
-  2. 2.
-    Differentiation between text tokens and graph tokens.
-It directs LLM to correctly distinguish between textual tokens and graph tokens within different segments of the prompt, thus preventing confusion.
-  3. 3.
-    Guidance for learning from multi-GNNs: It leads LLM to implicitly learn how to effectively combine strengths of different GNNs.
-Based on the above principles, we provide a prompt template used in our experiments.
-Details on the prompt template are given in Appendix D.
-Further, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously.
+The designed prompts include three major components: task instruction, node text and graph token.
+Specifically, task instruction consists of task specifications and also instructions that guide LLMs to ensemble multi-GNNs.
+Node text is used to inject textual attributes of target nodes and their neighbors.
+Due to the input length limitation, for each target node, we only consider up to 20 adjacent nodes.
+When the number of adjacent nodes for the target node exceeds 20 or when the total token count of the prompt surpasses the predefined limit, a random sampling of the adjacent nodes is performed.
+Further, graph tokens are derived from multi-GNNs, which contain rich graph structural information.
+After prompts are deigned, they are fed into LLMs to predict labels.
+Formally, we have:
+$o=LLM(x_{\text{instruction}},x_{\text{node-text}},\{x_{\text{GNN}_{i}}\}_{i=1}^{k}),$
+where $o$ is the predicted label, $x_{\text{instruction}}$ denotes the instruction text, $x_{\text{node-text}}$ is node textual descriptions and $x_{\text{GNN}_{i}}$ denotes the graph representation (graph token) generated by the $i$-th GNN.
+Details on the prompt template used in our experiments are given in Appendix D.
+Therefore, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously.
 In our experiments, we merge them into one step.
 5 EXPERIMENT
-Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods.
-LensGNN utilizes Baichuan2-13B-chat as the backbone LLM.
-We highlight the best score on each dataset in bold and underline the runner-up’s.
 Model type Model Cora PubMed ogbn-arXiv Citeseer Wiki-CS
@@
 DeBERTa 77.79 93.45 72.90 73.13 75.11
-LLM Ensembler LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
+SciBERT 83.21 95.26 73.11 77.74 76.83
+MedBERT 77.31 93.94 72.09 74.29 74.04
+LLM Ensembler
+LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
 LensGNN-[GCN+GIN] 91.88 94.47 75.89 76.64 83.47
@@
 LensGNN-ALL 90.40 95.68 75.91 79.31 81.78
+Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods.
+LensGNN utilizes Baichuan2-13B-chat as the backbone LLM.
+We highlight the best score on each dataset in bold and underline the runner-up’s.
+Here, LensGNN-ALL refers to LensGNN-[GCN+GAT+GIN].
 5.1 Experimental settings
@@
 Baselines.
-We compare our method with 19 baselines: MLP, GNN models, LM-based models and LLM-based models.
-Specifically, GNN models include GCN Kipf and Welling (2016), GAT Veličković et al.
-(2017), GIN Xu et al.
-(2019), GraphSAGE Hamilton et al.
+We compare our method with 23 baselines: MLP, GNN models, ensemble models, LM-based models and LLM-based models.
+Specifically, GNN models include GCN, GAT, GIN, GraphSAGE Hamilton et al.
 (2017a) and Graphormer Ying et al.
 (2021).
-LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019) and DeBERTa He et al.
-(2020).
+Ensemble models include Bagging Breiman (1996), Stacking Wolpert (1992) and AdaBoost Freund and Schapire (1997).
+LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019), DeBERTa He et al.
+(2020), SciBERT Beltagy et al.
+(2019) and MedBERT Feihong Yang (2021).
 LLM-based models include DGTL Qin et al.
@@
 (2024), LLaGA Chen et al.
+(2024), InstructGLM Ye et al.
 (2024), OFA Liu et al.
@@
 Setup.
-We use Baichuan2 Yang et al.
-(2023) and InternLM2.5 Cai et al.
+We use Baichuan2 and InternLM2.5 Cai et al.
 (2024) as the backbone LLM for LensGNN.
@@
 For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input.
+The ensemble experiments for more GNNs are presented in Appendix I.
+The hyperparameter analysis experiment is detailed in Appendix J.
+The analysis of input length is presented in Appendix K.
 For more details on experiment setup, refer to Appendix C.
+Detailed hyperparameter settings are provided in Appendix L.
+Additionally, in Appendix H, we include supplementary experiments utilizing LLama2-7B as the backbone.
 5.2 Node classification results
@@
 Table 3 evaluates the performance of LensGNN in the 20-shot setting.
+Table 4 provides a comparison between LensGNN and methods that ensemble GCN, GAT and GIN.
+We also experimented with traditional methods that ensemble LLM and GNNs.
+The experimental analysis is provided in Appendix G.
+Additionally, we conduct one-shot experiments, which are discussed in Appendix F.
 From these tables, we observe:
@@
 We speculate this is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data.
-Different from other LLM-based models, LensGNN integrates the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
-Table 2: Comparison on classification accuracy (%) with LLM-based models. * denotes the best variant of our model and - means that the results are missing from their original papers.
+Different from other LLM-based models, LensGNN ensembles the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
+(3) LensGNN leads standard ensemble methods across all datasets.
+This is because when ensembling multiple GNNs, standard ensemble methods can only leverage graph structural information.
+However, LLMs as ensembler have strong semantic understanding capability that further improves the model performance.
 Model Cora PubMed ogbn-arXiv
@@
 LensGNN*-Baichuan2-13B 91.88 95.68 75.91
-Table 3: Classification accuracy in 20-shot setting.
+Table 2: Comparison on classification accuracy (%) with LLM-based models.
+* denotes the best variant of our model and - means that the results are missing from their original papers.
+The best variant model is given in Appendix E.
 Model Cora PubMed
@@
 LensGNN*-InternLM2.5-7B 79.93 87.75
+Table 3: Classification accuracy (%) in 20-shot setting.
+Model Cora PubMed Citeseer WikiCS
+Bagging 87.16 82.67 77.35 76.97
+Stacking 86.25 84.85 78.08 74.12
+AdaBoost 84.91 83.36 76.09 75.41
+LensGNN-ALL 91.88 95.68 79.31 83.47
+Table 4: Comparison with ensemble methods (%).
 5.3 Graph classification results
-Table 4: Graph classification results.
-Model BACE BBBP ClinTox
+Model
+BACE BBBP ClinTox
 ACC AUC ACC AUC ACC AUC
@@
 LensGNN-ALL 80.59 80.33 88.53 85.37 98.99 97.63
+Table 5: Graph classification results.
 We further evaluate the effectiveness of LensGNN on the graph classification task.
 We use classification accuracy (ACC) and AUC scores as the evaluation metrics.
-The results are given in Table  4.
+The results are given in Table  5.
 From the table, LensGNN outperforms GNN baselines in most cases, which shows that combining GNNs and LLMs is inspiring.
@@
 This illustrates that by integrating the strengths of various GNNs, LensGNN can effectively enhance the overall performance.
+On Cora and Wiki-CS, the small training sets introduce more noise during training.
+Since our model relies on an LLM as backbone—and LLMs typically require large-scale data for effective learning—we attribute the underperformance of LensGNN-ALL compared to LensGNN-[GCN+GIN] primarily to the limited training data and elevated noise levels.
+These factors hinder the LLM from adequately learning to align all three GNNs (GCN, GAT, and GIN) simultaneously.
 Impact of Alignment.
@@
 Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, thereby improving classification performance.
-Table 5: Ablation study results.
 GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
@@
 GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
+Table 6: Ablation study results.
 5.5 Performance with different backbones
-Table 6: The classification accuracy (%) with different LLMs on Cora and PubMed.
 Model Bert-base T5-base (Encoder only) T5-base Falcon-7B InternLM2.5-7B-chat Baichuan2-13B
@@
 Ensemble All 94.17 93.51 93.95 94.97 95.13 95.68
+Table 7: The classification accuracy (%) with different LLMs on Cora and PubMed.
 To study the impact of backbone models on LensGNN, we use different backbones, including three small LMs: BERT-base Devlin (2018), T5-base Raffel et al.
@@
 (1) Small LMs BERT-base, T5-base, and T5-base (Encoder only) perform well when applied to a single GNN, but the performance drops significantly when ensembling multiple GNN models.
-This suggests that models with less parameters lack the capability to integrate multiple GNNs.
+This suggests that models with less parameters lack the capability to ensemble multiple GNNs.
 (2) For LLMs, they generally perform better than LM backbones.
@@
 LensGNN adopts two phases of alignment: the multi-GNN alignment aims to map node representations from different GNNs into the same low-dimensional space, while the GNN-LLM alignment injects graph representation as graph tokens into LLM.
-After that, we perform supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology.
+After that, we performed supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology.
 Finally, we conducted extensive experiments to show the effectiveness of our ensembling model LensGNN.
-In particular, our experimental results showed that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
+In particular, the results show that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
 7 LIMITATIONS
@@
 For Wiki-CS, we split the dataset as suggested in Mernyei and Cangea (2022).
-Table 7: Datasets statistics.
 Dataset #Graphs Avg.# Nodes Avg.# Edges # Classes
@@
 ClinTox 1,491 26.1 55.5 2
+Table 8: Datasets statistics.
 Appendix B Baselines
@@
 Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data.
+(3) Ensemble Models: Bagging Breiman (1996) is an ensemble method that improves model stability and accuracy by training multiple base learners on bootstrap samples and aggregating their predictions through voting or averaging.
+Stacking Wolpert (1992) is an ensemble technique that combines predictions from multiple base learners to train a meta-learner, enhancing overall predictive performance through stacked generalization.
+AdaBoost Freund and Schapire (1997) is an adaptive boosting algorithm that sequentially trains weak learners, adjusts sample weights, and combines their predictions to create a strong classifier with improved accuracy.
 (4) LM Based Models: BERT Devlin (2018) is a groundbreaking model in natural language processing.
@@
 (2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
+SciBERT Beltagy et al.
+(2019) is a BERT-based language model pre-trained on a large multi-domain corpus of scientific publications to achieve state-of-the-art performance on various scientific NLP tasks.
+MedBERT Feihong Yang (2021) is a series of BERT and ALBERT models pre-trained on a large Chinese clinical corpus, which outperforms baseline models on named entity recognition and text classification tasks within the Chinese clinical NLP domain.
 Figure 5: A prompt template for node classification task.
@@
 This innovative approach enables LLMs to comprehend and interpret the structural components of graphs, thereby demonstrating superior generalization in both supervised and zero-shot graph learning tasks.
+Additionally, GraphGPT employs Baichuan-7B and Vicuna-7B as its backbone models.
 LLaGA Chen et al.
@@
 LLaGA achieves this by reorganizing graph nodes into structure-aware sequences and mapping them into the token embedding space through a versatile projector, demonstrating superior versatility, generalizability, and interpretability across various datasets and tasks.
+LLaGA employs Vicuna-7B as its backbone model.
 InstructGLM Ye et al.
 (2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning.
+InstructGLM employs Llama2-7B as its backbone model.
 DGTL Qin et al.
-(2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs.
+(2024) is an LLM-first-GNN-later method, which means first using LLMs to process text before training GNNs.
+DGTL incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs.
+DGTL employs Llama2-13B as its backbone model.
 SNS-GPT4 Li et al.
@@
 It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily.
+SNS-GPT4 employs GPT3.5-turbo and GPT4 as its backbone models.
+Notably, SNS-GPT4 demonstrates better performance than TAPE He et al.
+(2023) (another LLM-first-GNN-later approach) in its reported results.
 GAugLLM Fang et al.
@@
 It leverages advanced large language models like Mistral to enhance self-supervised graph learning.
+GAugLLM employs Mistral 8*7b, ChatGPT-3.5 and Llama2-13b as its backbone models.
 Baichuan2 Yang et al.
@@
 OFA achieves this by converting graph data into text-attributed graphs (TAGs), using language models to encode diverse text attributes into a common embedding space, and introducing a novel graph prompting paradigm for in-context learning without fine-tuning.
+OFA employs Llama2 as its backbone model.
 GOFA Kong et al.
@@
 It achieves this by interleaving GNN layers into a frozen pre-trained LLM, allowing for self-supervised pretraining on graph data and fluidity in handling various graph-related tasks.
-Table 8: Hyperparameter Search Spaces.
+OFA employs Llama2 and Mistral as its backbone models.
+Appendix C Experiment setup
+We implement LensGNN by PyTorch.
+We primarily use Baichuan2 Yang et al.
+(2023) and InternLM2.5 Cai et al.
+(2024) as the backbone LLM for LensGNN.
+In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers.
+We perform grid search to fine-tune hyperparameters based on the validation set.
+Details on the search space is given in Table 9.
+We utilize LoRA+ Hayou et al.
+(2024) for fine-tuning.
+Some fixed LoRA settings include capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step.
+We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1.
+For the training of LLMs, we utilize Llama-Factory Zheng et al.
+(2024) as a framework.
+For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input.
+For baselines that report results on the adopted datasets, we directly report the results from their original papers.
+For those whose results are missing, we leave them blank.
+For Baichuan2, we directly use it without fine-tuning.
+We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
+In addition, in the era of LLMs, due to the high cost, it is difficult to run experiments for baselines.
+For fair comparison, a widely adopted approach is to make comparison on the results for baselines reported from their original papers because their authors have well fine-tuned these models.
+This explains why only part of results are included in Table 2 and why we did not reimplement some baselines.
 Hyperparameter Search space
@@
 Graph token $t$ $\{1,2,4,8,16\}$
-Table 9: Detailed hyperparameters of aligning multi-GNNs.
-Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
+Table 9: Hyperparameter Search Spaces.
+Appendix D Prompt design
+The prompt design follows three key principles:
+  1. 1.
+    Simultaneous input of textual strings from target node and its adjacent neighbors, graph tokens of the node, and task specification.
+Existing studies Kipf and Welling (2016); Veličković et al.
+(2017); Verma et al.
+(2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction.
+Further, clear task instructions are necessary for accurate prediction.
+  2. 2.
+    Differentiation between text tokens and graph tokens.
+It directs LLM to correctly distinguish between textual tokens and graph tokens within different segments of the prompt, thus preventing confusion.
+  3. 3.
+    Guidance for learning from multi-GNNs: It leads LLM to implicitly learn how to effectively combine strengths of different GNNs.
+The prompt is divided into two parts: Instruction and Input.
+The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content.
+The latter provides the specific text of target node or graph and their multiple GNN representations.
+An example of the prompt template can be seen in Figure 5. For graph classification, we follow MolXPT Liu et al. (2023) to design prompts for molecular graphs as shown in Figure 6.
+Appendix E Best variants
+The best variants in Table 2 and Table 3 are listed in Table 10 and Table 11, respectively.
+Backbone LLM Cora PubMed ogbn-arXiv
+InternLM2.5-7B LensGNN-ALL LensGNN-ALL LensGNN-[GAT+GIN]
+Baichuan2-7B LensGNN-[GCN+GAT] LensGNN-ALL LensGNN-ALL
+Baichuan2-13B LensGNN-[GCN+GIN] LensGNN-ALL LensGNN-ALL
+Table 10: Best variants in Table 2.
+Backbone LLM Cora PubMed
+Baichuan2-7B LensGNN-ALL LensGNN-ALL
+InternLM2.5-7B LensGNN-ALL LensGNN-[GAT+GIN]
+Table 11: Best variants in Table 3.
+Appendix F One-shot results
+We further conduct experiments on one-shot setting, the results are presented in Table 12. From the table, we see that
+Model Cora PubMed Citeseer Wiki-CS
+GCN 0.1890 0.4155 0.3242 0.2873
+GAT 0.2543 0.4993 0.3506 0.4094
+GIN 0.2021 0.4071 0.2763 0.3278
+LensGNN 0.5938 0.8379 0.5332 0.3047
+Table 12: Classification accuracy in one-shot setting.
+our method still achieves outstanding performance on the Cora, PubMed, and Citeseer datasets, which shows the strong generalizability.
+For WikiCS, due to the very scare labeled data, GNN ensembling is adversely affected.
+Appendix G Traditional methods ensembling LLM and GNNs
+We have conducted additional experiments employing traditional ensemble learning methods to ensemble the predictions from both the LLM and GNNs (GCN, GAT and GIN) on 4 datasets.
+The results are as Table 13:
+Method LLM Cora PubMed Citeseer Wiki-CS
+Bagging Baichuan2-13B 86.90 86.16 75.55 76.47
+Stacking Baichuan2-13B 85.24 85.70 70.53 67.35
+LensGNN* Baichuan2-13B 91.88 95.68 79.31 83.47
+Table 13: Comparison with traditional ensemble of LLM and GNNs. (%)
+As demonstrated in our experiments, LensGNN consistently outperforms all conventional ensemble methods.
+Appendix H LLama2-7B as backbone
+We have supplemented our experiments by implementing LLaMA2-7B as the backbone LLM.
+We conducted these experiments on a medium-scale dataset, PubMed, as a representative, with results presented in Table 14:
+Model PubMed
+MLP 82.40
+GCN 83.78
+GAT 83.59
+GIN 82.03
+GraphSAGE 87.01
+Graphormer 88.75
+BERT 93.91
+SentenceBERT 92.49
+DeBERTa 93.45
+GraphGPT-MIX-7B 64.76
+LLaGA-HO-7B(GENERAL) 75.01
+InstructGLM-Llama-7B 93.84
+DGTL 87.10
+SNS-GPT4 93.80
+GAugLLM 74.15
+Baichuan2-13B 36.04
+LensGNN*-LLaMA2-7B 94.59
+Table 14: Performance comparison (%) on PubMed dataset using LLaMA2-7B as backbone. (%)
+Experimental results demonstrate that our model with LLaMA2-7B as the backbone also achieves superior accuracy compared to: (1) GNN baselines, (2) LM-based methods, and (3) other LLM-based approaches.
+Appendix I Ensemble more GNN
+We further incorporate a larger number of GNNs to evaluate how LensGNN scales and whether the performance continues to improve.
+The results are presented in Table 15.
+The experiment is based on the backbone of Internlm2.5-7B-chat, with the number of graph tokens $t=4$.
+Overall, we see that with the increase in the number of GNNs, the model performance improves further.
+However, from the results on Cora and Pubmed, simply adding GNNs will not bring sustained improvement.
+We speculate that this is because more GNNs will introduce more graph tokens and also noise.
+The increased input length will bring difficulty in model learning, so does noise.
+Based on the empirical findings presented in Table 15, we propose a heuristic recommendation for researchers to select 3-4 high-performing GNNs (e.g., GCN, GAT, GIN) when constructing ensembles for new datasets.
+Notably, the inclusion of additional GNN models incurs minimal computational overhead, allowing for flexible expansion until validation accuracy approaches saturation—the point at which further performance improvements become negligible.
+This strategy balances model diversity and efficiency while maximizing predictive accuracy.
+Variants of LensGNN Number of GNNs Cora PubMed Citeseer Wiki-CS
+LensGNN-GCN 1 0.9059 0.9490 0.7836 0.8294
+LensGNN-[GCN+GAT] 2 0.9003 0.9503 0.7868 0.8306
+LensGNN-[GCN+GAT+GIN] 3 0.9114 0.9508 0.7884 0.8334
+LensGNN-[GCN+GAT+GIN+GraphSAGE] 4 0.9095 0.9513 0.7852 0.8332
+LensGNN-[GCN+GAT+GIN+APPNP] 4 0.9040 0.9500 0.7915 0.8366
+LensGNN-[GCN+GAT+GIN+GraphSAGE+APPNP] 5 0.9022 0.9472 0.7931 0.8417
+Table 15: Classification accuracy when integrating more GNNs.
+Appendix J Additional hyperparameter sensitivity analysis
+We further conduct hyperparameter sensitivity analysis on Citeseer and Wiki-CS, whose results are given in Figure 7.
+From the figure, we see that the model can also achieve good results at $t=8$ in terms of the overall performance, which is similar as the results in Figure 4.
+Figure 7: Hyperparameter sensitivity analysis.
+Appendix K Input length analysis
+We investigate the impact of LLM input length on model efficiency and performance across Cora, Citeseer, and Wiki-CS.
+Specifically, we established different limits on total number of input tokens.
+When exceeded, some neighbor text is randomly truncated.
+Within this constraint, more text introduces additional noise but also provides richer useful information, leading to better model performance—albeit at increased computational cost.
+The experimental results are presented in Figure 8.
+Figure 8: Input length analysis.
+Our experiments using Baichuan-7B as the backbone record sample statistics with total token length limits ranging from 0.5K to 1.75K, including average input string length, training time, and test accuracy.
+While Cora’s shorter texts show negligible variations, both Citeseer and Wiki-CS exhibit a clear trend: longer inputs increased training time but also improved accuracy.
+Appendix L Detailed hyperparameter settings
+The details on the setting of hyperparameters in the first and second stages of LensGNN are shown in Tables 16 and 17, respectively.
+The hyperparameters for the 20-shot experiment are shown in Tables 18 and 19.
+Model Dataset Hidden size
+GNN representation
+size
+Number of
+layers
+Learning rate of
+GNN encoder
+Initial learning rate of
+the classifier
+Dropout Rate Weight Decay
 GCN Cora 256 40960 2 0.0002 0.0002 0.25 0.001
@@
 ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
-Table 10: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
-Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
+Table 16: Detailed hyperparameters of aligning multi-GNNs.
+Model Dataset
+Lora dropout
+Loraplus
+lr ratio
+Training
+batch size
+Learning rate
+Early stop
+epoch
+Warmup ratio
 Baichuan2-13B Cora 0.1 16 4 0.0001 3 0.1
@@
 Pubmed 0.1 16 4 0.0001 2 0.1
-Appendix C Experiment setup
-We implement LensGNN by PyTorch.
-We primarily use Baichuan2 Yang et al.
-(2023) and InternLM2.5 Cai et al.
-(2024) as the backbone LLM for LensGNN.
-In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers.
-We perform grid search to fine-tune hyperparameters based on the validation set.
-Details on the search space is given in Table 8.
-We utilize LoRA+ Hayou et al.
-(2024) for fine-tuning.
-Some fixed LoRA settings include capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step.
-We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1.
-For the training of LLMs, we utilize LLaMA-Factory Zheng et al.
-(2024) as a framework.
-For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input.
-For baselines that report results on the adopted datasets, we directly report the results from their original papers.
-For those whose results are missing, we leave them blank.
-For Baichuan2, we directly use it without fine-tuning.
-We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
-Appendix D Prompt design
-The prompt is divided into two parts: Instruction and Input.
-The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content.
-The latter provides the specific text of target node or graph and their multiple GNN representations.
-An example of a prompt template can be seen in Figure 5.
-For graph classification, we follow MolXPT Liu et al.
-(2023) to design prompts for molecular graphs as shown in Figure 6.
-Appendix E Detailed hyperparameter settings
-The details on the setting of hyperparameters in the first and second stages of LensGNN  are shown in Tables 9 and 10, respectively.
-The hyperparameters for the 20-shot experiment are shown in Tables 11 and 12.
-Table 11: Detailed hyperparameters of aligning multi-GNNs (20-shot).
-Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
+Table 17: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
+Model Dataset Hidden size
+GNN representation
+size
+Number of
+layers
+Learning rate of
+GNN encoder
+Initial learning rate of
+the classifier
+Dropout Rate Weight Decay
 GCN Cora 256 32768 2 0.0002 0.0005 0.25 0.001
@@
 Pubmed 256 32768 2 0.0001 0.0005 0.25 0.001
-Table 12: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune) (20-shot).
-Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
+Table 18: Detailed hyperparameters of aligning multi-GNNs (20-shot).
+Model Dataset
+Lora dropout
+Loraplus
+lr ratio
+Training
+batch size
+Learning rate
+Early stop
+epoch
+Warmup ratio
 Baichuan2-7B Cora 0.15 16 4 0.0001 8 0.1
@@
 Pubmed 0.05 16 4 0.0001 9 0.1
+Table 19: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune) (20-shot).